### v1

# On the Jacobs-de Leeuw-Glicksberg decomposition

Abstract
For any JdLG-admissible representation $\pi$ of a semigroup $S$, we show that the reversible part is weakly equivalent to a unitary representation on a Hilbert space that decomposes into a direct sum of finite dimensional representations, and we give an alternative characterization of the almost weakly stable part in terms of the unique invariant mean on the space of weakly almost periodic functions. In the case that $S$ is a bi-amenable measured semigroup, we characterize the almost weakly stable part using invariant means and averages along Følner sequences. Moreover, we give a description of the unique projection onto the reversible part whose kernel is the almost weakly stable part in terms of ultrafilters.
1 Introduction
The compact-weak mixing decomposition states that for any unitary operator $U$ acting on a Hilbert space $\mathcal{H}$, we have a decomposition $\mathcal{H}=\mathcal{H}_{c}\oplus\mathcal{H}_{w}$ with
[…]
$\mathcal{H}_{w}$ $=\left\{\xi\in\mathcal{H}\ |\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\xi\rangle|=0\right\}=\left\{\xi\in\mathcal{H}\ |\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\eta\rangle|=0\ \forall\ \eta\in\mathcal{H}\right\}.$
Koopman and von Neumann [33] were the first to show that a dynamical system is weakly mixing if and only if it has no nontrivial eigenfunctions. A version of the compact-weak mixing decomposition for locally compact groups was found by Godement [27] (see also [40, Appendix 10.D]), and a version for amenable semigroups was found by Dye [18]. In between Godement and Dye, the compact-weak mixing decomposition was extended to various semigroups of operators on a Banach space by Jacobs, de Leeuw, and Glicksberg [30, 17], and is now known as the Jacobs-de Leeuw-Glicksberg (JdLG) decompoisition. Several applications of the JdLG-decomposition can be found in the literature, e.g. it is used in ergodic Ramsey theory [7, Theorem 2.3], the proof of the Erdös sumset conjecture[36], the Fürstenberg-Zimmer structure theorem [19], a proof of the Wiener-Wintner theorem [1, Proposition 2.2], and more generally in the study of weighted ergodic theorems [24, 4]. Further uses include results related to the entangled ergodic theorem [23], the Perron-Frobenius theory for completely positive operators on $W^{*}$-algebras [5], and evolution equations [25, Section V.2].
A modern treatment of the JdLG decomposition is given in [21, Chapter 16]. In the present paper we will have a heavy focus on representations of semigroups, so we give a statement of the JdLG decomposition as follows.
Theorem 1.1 (cf. [13, Theorem 6.2.18] or [28, Theorem 3.8.5]).
Let $S$ be a semitopological semigroup which admits a bi-invariant mean $m$ on $W(S)$, the space of weakly almost periodic functions on $S$. Let $E$ be a Banach space and let $\pi$ be a relatively weakly compact representation of $S$ on $E$. For $\mathscr{S}=c\ell_{\sigma}(\pi(S))$, the weak closure of $\pi(S)$, we have $E=E_{r}\oplus E_{\text{aws}}$, where
$E_{r}$ $=\{\xi\in E\ |\ \forall\ u\in\mathscr{S},\exists\ v\in\mathscr{S}\text{ with }vu\xi=\xi\}\text{ and}$
$E_{\text{aws}}$ $=\{\xi\in E\ |\ u\xi=0\text{ for some }u\in\mathscr{S}\},$
and both of $E_{r}$ and $E_{\text{aws}}$ are closed $\mathscr{S}$-invariant subspaces.
A semigroup $\mathscr{T}$ of operators on a Banach space $E$ is JdLG-admissible if it is relatively weakly compact and we have a decomposition $E=E_{r}\oplus E_{\text{aws}}$ for the weak closure $\mathscr{S}$ of $\mathscr{T}$. $E_{r}$ is called the reversible part of $E$ and $E_{\text{aws}}$ is the almost weakly stable part of $E$. We now recall the list of examples of JdLG-admissible semigroups.
Theorem 1.2 (cf. [21, Chapter 16.3]).
Let $E$ be a Banach space and $\mathscr{T}$ a semigroup of bounded linear operators on $E$. Then $\mathscr{T}$ is JdLG-admissible in the following cases:
  1. (i)
    $\mathscr{T}$ is abelian and relatively weakly compact.
  2. (ii)
    $E$ is a Hilbert space and $\mathscr{T}$ consists of contractions.
  3. (iii)
    $E=L^{1}(X,\mu)$ with $(X,\mu)$ a probability space, and $\mathscr{T}$ consists of Markov operators.
  4. (iv)
    $E$ and its dual space $E^{\prime}$ both are strictly convex, and $\mathscr{T}$ is relatively weakly compact.
  5. (v)
    $E=L^{p}(X,\mu)$ for some measure space $(X,\mu)$ and $1<p<\infty$, and $\mathscr{T}$ consists of contractions.
We call a representation $\pi$ of a semigroup $S$ on a Banach space $E$ JdLG-admissible if $\mathscr{T}:=(\pi_{s})_{s\in S}$ is JdLG-admissible. It is known that a relatively weakly compact semigroup $\mathscr{T}$ of operators is JdLG-admissible if and only if there exists a bi-invariant mean $M$ on $C_{b}(\mathscr{T})$ (cf. [17, Theorem 4.11]). If $S$ is a locally compact group, an amenable group, or a bi-amenable measured semigroup, then it has a unique invariant mean on $W(G)$. Consequently, the following result follows immediately from Theorem 1.1.
Corollary 1.3.
Let $S$ be either a locally compact group, an amenable group, or a bi-amenable measured semigroup. If is $E$ a Banach space and $\pi$ is a relatively weakly compact representation of $S$ on $E$, then $\pi$ is JdLG-admissible. In particular, if $\mathscr{T}$ is a semigroup of bounded linear operators on $E$ with $c\ell_{\sigma}(\mathscr{T})=c\ell_{\sigma}(\pi(S))$, then $\mathscr{T}$ is JdLG-admissible.
We see that Corollary 1.3 is a generalization of Theorem 1.2(i) by taking $\mathscr{S}=S$, $\pi=\text{Id}$, and recalling that every discrete abelian semigroup endowed with the counting measure is bi-amenable.
The main goal of the present paper is to give alternative characterizations of the reversible and almost weakly stable parts of the JdLG decomposition. Results relating the compact-weak mixing decomposition to the algebra of ultrafilters on $\mathbb{N}$ have been investigated in [8, 9, 12, 35], and in Section 5 we show that many of these results extend to JdLG-admissible semigroups of operators on a Banach space. Alternative characterizations of $E_{r}$ are already given in Proposition 16.29 and Theorem 16.31 of [21], and we will build upon these characterizations in Section 3. Alternative characterizations of $E_{\text{aws}}$ in terms of flight functions were already alluded to in [17, Corollary 5.9], [18, Theorem 1], and [28, Page 88], but were not stated explicitly prior to this paper. In particular, the next result follows from Theorem 3.4 and Lemma 4.1.
Theorem 1.4.
Let $E$ be a Banach space, $\mathscr{T}\subseteq\mathcal{L}(E)$ a JdLG-admissible semigroup of operators, and $E=E_{r}\oplus E_{\text{aws}}$ the JdLG-decomposition.
  1. (i)
    The restriction of $\mathscr{T}$ to $E_{r}$ is weakly equivalent to a unitary representation $U$ of $\mathscr{T}$ on a Hilbert space $\mathcal{H}$. Furthermore, the representation $U$ decomposes into a direct sum of finite dimensional representations.
  2. (ii)
    Let $M$ be a bi-invariant mean on $C_{b}(\mathscr{T})$. We have
    $E_{\text{aws}}=\left\{\xi\in E\ |\ \forall\ f^{\prime}\in E^{\prime},M\left(t\mapsto\left|f^{\prime}(\pi_{t}\xi)\right|\right)=0\right\}.$ (1)
When $S=\mathbb{N}$ in Theorem 1.1, alternative characterizations of $E_{\text{aws}}$ resembling the characterizations of $\mathcal{H}_{w}$ are stated in [21, Theorem 16.34]. In Theorem 4.2 we extend these alternative characterizations to various classes of amenable semigroups. For now, we only state a particularly nice corollary of Corollary 1.3, Corollary 3.2, and Theorem 4.2.
Theorem 1.5.
Let $G$ be a locally compact amenable group, $\lambda$ a left-Haar measure, $\mathcal{F}=(F_{i})_{i\in I}$ a left-Følner net, $E$ a Banach space, and $\pi$ a relatively weakly compact representation of $G$ on $E$. Then $\pi$ is JdLG-admissible, and we further have that
$E_{r}$ $=\{\xi\in E:\pi(G)\xi\text{ is relatively compact in }E\}\text{, and}$
$E_{aws}$ $=\left\{\xi\in E:\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{g}\xi,x^{\prime}\right\rangle\right|d\lambda(g)=0\text{ for all }x^{\prime}\in E^{\prime}\right\}$
Remark.
In many cases of interest, the Banach space $E$ is reflexive. Consequently, it is worth recalling that a set of bounded linear operators on a reflexive Banach space $E$ is relatively weakly compact if and only if it is uniformly bounded.
The structure of this paper is as follows. In Section 2, we collect known facts about topological semigroups, weak almost periodicity, JdLG-admissible semigroups, amenability, representations of semigroups, and ultrafilters. In Section 3 we study the reversible part of the JdLG-decomposition. In Section 4 we study the almost weakly stable part of the JdLG-decomposition. In Section 5 we study the relationship between the JdLG-decomposition and the algebra of the Stone-Čech compactification.
Acknowledgements: The second author acknowledges being supported by grant 2019/34/E/ST1/00082 for the project “Set theoretic methods in dynamics and number theory,” NCN (The National Science Centre of Poland), and the first author was supported by the same grant for a 2 research visits. The second author would like to thank Uta Freiberg for funding a research visit to TU Chemnitz. We would like to thank Yulia Kuznetsova for providing us with Example 4.3. We would also like to thank Rigoberto Zelada for helpful discussions about Theorem 5.2.
2 Preliminaries
2.1 Topological semigroups and weak almost periodicity
A semigroup $(S,\cdot)$ is a set $S$ with an associative binary operatation $\cdot:S\times S\rightarrow S$. If $S$ possesses an identity element $e$, then $S$ is called a monoid. If $S$ is also a Hausdorff topological space and the map $\cdot$ is continuous, then $S$ is a topological semigroup [^4] We warn the reader that in [17], the term topological semigroup refers to what we call a semitopological semigroup.. If $S$ is a Hausdorff topological space for which $\cdot$ is separately continuous, i.e., for each $s\in S$ the maps $L_{s}(t):=s\cdot t$ and $R_{s}(t):=t\cdot s$ are continuous, then $S$ is a semitopological semigroup. For the rest of this paper all semigroups are assumed to be semitopological unless it is explicitly stated otherwise. By abuse of notation, we let $R$ act on $C_{b}(S)$ and $L^{\infty}(S,\lambda)$ for some Borel measure $\lambda$, by $R_{s}(f)(t)=f(R_{s}(t))=f(t\cdot s)$, and similarly for $L$. A function $f\in C_{b}(S)$ is weakly almost periodic if $\{R_{s}f\ |\ s\in S\}$ is relatively weakly compact. We let $W(S)$ denote the space of weakly almost periodic functions on $S$. De Leeuw and Glicksberg [17, Theorem 2.7] showed that if $S$ is a compact semigroup, then $W(S)=C(S)$. For $X\in\{C_{b}(S),W(S),L^{\infty}(S,\lambda)\}$, a left-invariant mean on $X$ is a positive linear functional $m:X\rightarrow\mathbb{C}$ of norm $1$ that satisfies $m(L_{s}f)=m(f)$ for all $s\in S$ and all $f\in X$, and a right-invariant mean is defined similarly. The semigroup $S$ admits a bi-invariant mean on $X$ if there exists a left-invariant mean $m$ on $X$ that is also a right-invariant mean.
In [17, Theorem 5.8] it is shown that if $W(S)$ admits a bi-invariant mean, then $W(S)$ admits a unique bi-invariant mean. It is also known that $W(S)$ admits a bi-invariant mean if and only it it admits a left-invariant mean and a (potentially different) right-invariant mean. Now let $S$ be a semigroup with bi-invariant mean $m$ on $W(S)$. If $\mathscr{S}$ is a compact semigroup and $\pi:S\rightarrow\mathscr{S}$ is a continuous homomorphism with dense image, then $\tilde{\pi}:C(\mathscr{S})\rightarrow W(S)$ is given by $(\tilde{\pi}f)(s)=f(\pi_{s})$ is well defined and $M(f):=m(\tilde{\pi}f)$ is the bi-invariant mean on $W(\mathscr{S})$ (see [17, Lemma 5.2]). There exists a compact semigroup $S^{w}$, the weakly almost periodic compactification of $S$, and a continuous injective homomorphism $i:S\rightarrow S^{w}$ with dense image, such that for any $f\in W(S)$ there is a unique $f^{w}\in W(S^{w})$ satisfying $f^{w}(i(s))=f(s)$ for all $s\in S$ (cf. [17, Theorem 5.3]). By abuse of notation, we let $i:W(S)\rightarrow W(S^{w})$ denote the map $f\mapsto f^{w}$, and we observe that $i$ is an algebra isomorphism. It is known that $W(S^{w})$ admits a bi-invariant mean $M$ if and only if $K(S^{w})$, the smallest two-sided ideal of $S^{w}$, is a compact topological group and $M(f)=\int_{K(S^{w})}fd\mu$, where $\mu$ is the normalized Haar measure on $K(S^{w})$.
Now suppose that $W(S^{w})$ admits a bi-invariant mean $M$ and a left-invariant mean $M_{\ell}$. For $f\in W(S^{w})$ and $g\in K(S^{w})$, we have $M_{\ell}(f)=M_{\ell}(L_{g}f)$, hence $M_{\ell}(f)$ is determined by the values of $f$ on $K(S^{w})$. Tietze’s Extension Theorem tells us that any $f\in C(K(S^{w}))$ extends to a member of $W(S^{w})$, so the left-invariant means on $W(S^{w})$ are in a one-to-one correspondence with the left-invariant means on $C(K(S^{w}))$. The uniqueness of the Haar measure tells us that there is exactly one left-invariant mean on $C(K(S^{w}))$, hence $M_{\ell}=M$. Since the means on $W(S)$ are in a one-to-one correspondence with the means on $W(S^{w})$, the analogous results holds for $W(S)$.
A group $(G,\cdot)$ is a topological group if it is a topological semigroup and the inversion map is a continuous map. All groups that we consider in this paper will be topological groups. De Leeuw and Glicksberg [17, Theorem 5.14] investigated which topological groups $G$ admit a bi-invariant mean on $W(G)$. 66 years later, Ryll-Nardzewski [39] proved a fixed point theorem that allows one to deduce that all locally compact groups $G$ admit a bi-invariant mean on $W(G)$. For further discussions on this topic the reader is referred to [28, 13, 14].
2.2 JdLG-admissible semigroups
For a Banach space $E$ we let $\mathcal{L}(E)$ denote the space of bounded linear operators on $E$. $\mathcal{L}(E)$ is naturally a semitopological semigroup when endowed with the operation of composition as well as the weak operator topology. In [21, Chapter 16.3], a semigroup $\mathscr{T}\subseteq\mathcal{L}(E)$ is called JdLG-admissible if it is relatively compact and $\mathscr{S}:=\operatorname{cl}_{\sigma}(\mathscr{T})$ contains a unique minimal idempotent $Q$. We will now show that this definition is equivalent to our definition. To this end, let us temporarily refer to the definition of JdLG-admissible from Section 1 as JdLG-decomposable in order to avoid confusion. If $\mathscr{T}$ is a relatively weakly compact semigroup, then [21, Theorem 16.5] says that the following are equivalent:
  1. 1.
    $\mathscr{T}$ is JdLG-admissible.
  2. 2.
    $K(\mathscr{S})$, the smallest two-sided ideal of $\mathscr{S}$, is a group.
  3. 3.
    $\mathscr{S}$ has a unique minimal right ideal and a unique minimal left ideal.
[21, Theorem 16.3] tells us that every minimal right ideal of $\mathscr{S}$ is closed, and [21, Lemma 16.4] tells us that $K(\mathscr{S})$ is the union of all minimal right ideals of $\mathscr{S}$. It follows that when $\mathscr{T}$ is JdLG-admissible, $K(\mathscr{S})$ is a closed (hence compact) subsemigroup of $\mathscr{S}$. [17, Theorem 2.1] says that a compact semitopological semigroup that is algebraically a group must be a topological group. It follows that $\mathscr{T}$ is JdLG-admissible if and only if $K(\mathscr{S})$ is a compact topological group. [17, Theorem 4.11] states that $K(\mathscr{S})$ is a compact topological group if and only if $\mathscr{T}$ is JdLG-decomposable. Putting everything together, we see that $\mathscr{T}$ is JdLG-decomposable if and only if it is JdLG-admissible.
2.3 Amenable semigroups and Følner nets
The pair $(S,\lambda)$ is a measured semigroup if $S$ is a semitopological semigroup endowed with a Borel measure $\lambda$. The main cases of interest are when $S$ is a discrete semigroup and $\lambda$ is the counting measure, or when $S$ is a subsemigroup of a locally compact topological group and $\lambda$ is the restriction of a Haar measure. The pair $(S,\lambda)$ is left-amenable if there exists a left-invariant mean $m$ on $L^{\infty}(S,\lambda)$, and it is bi-amenable if there exists a bi-invariant mean on $L^{\infty}(S,\lambda)$. For the rest of this section let $S$ denote a locally compact semitopological semigroup and let $\lambda$ be a Radon measure on $S$. A left-Følner net is a net $\mathcal{F}=(F_{i})_{i\in I}$ is a net of compact subsets of $S$ satisfying
$\lim_{i}\frac{\lambda(sF_{i}\triangle F_{i})}{\lambda(F_{i})}=0\text{ for all }s\in S.$ (2)
$\mathcal{F}$ is a two-sided Følner net if it is sumltaneously a left-Følner net and a right-Følner net. If $S$ possesses a left/right/two-sided Følner net $\mathcal{F}=(F_{i})_{i\in I}$ then $S$ is left/right/bi amenable. To see this, we define a net of functionals $(\mu_{i})_{i\in I}$ on $L^{\infty}(S,\lambda)^{\prime}$ by $\mu_{i}(f)=\frac{1}{\lambda(F_{i})}\int_{F_{i}}f(s)d\lambda(s)$ for all $f\in L^{\infty}(S,\lambda)$. Since the unit ball of $L^{\infty}(S,\lambda)^{\prime}$ is weak compact, we let $\mu$ be any cluster point of $(\mu_{i})_{i=1}^{\infty}$ and we see that $\mu$ is a left/right/two-sided invariant mean on $L^{\infty}(S,\lambda)$. We say that $\mu$ is a mean generated by $\mathcal{F}$ and we denote the set of such means by $\mathcal{M}(\mathcal{F})$.
There are examples of discrete semigroups that are amenable but do not possess a Følner net. However, if $S$ is a discrete cancellative semigroup, then $S$ is amenable if and only if it possesses a Følner net. It is also well known that if $S$ is a discrete abelian semigroup, then $S$ possesses a two-sided Følner net. We refer the reader to [37, Chapter 4.22] for further discussion on Følner conditions and amenability for discrete semigroups.
Now suppose that $S$ is a left-amenable semigroup. We define the upper (left) and lower (left) Banach densities on $S$ by
$d^{*}(A)$ $=\sup\{m(A)|m\text{ is a left-invariant mean on }S\},\text{ and}$
$d_{*}(A)$ $=\inf\{m(A)|m\text{ is a left-invariant mean on }S\}=1-d^{*}(A^{c}).$
If we further assume that $S$ possesses a left-Følner net $\mathcal{F}=(F_{i})_{i\in I}$, then we define the upper and lower densities along $\mathcal{F}$ by
$\overline{d}_{\mathcal{F}}(A):=\limsup_{i}\frac{\lambda(A\cap F_{i})}{\lambda(F_{i})}\text{ and }\underline{d}_{\mathcal{F}}(A):=\liminf_{i}\frac{\lambda(A\cap F_{i})}{\lambda(F_{i})},$
respectively. If $\overline{d}_{\mathcal{F}}(A)=\underline{d}_{\mathcal{F}}(A)$, then we denote the common value by $d_{\mathcal{F}}(A)$. We define the upper (left) and lower (left) Følner densities on $S$ by
[…]
$\underline{d}(A)$ $=\inf\left\{\overline{d}_{\mathcal{F}}(A)|\mathcal{F}\text{ is a (left) f\o lner net in }S\right\}=1-\overline{d}(A^{c}).$
If $S$ is a discrete semigroup and $\lambda$ is the counting measure, then [26, Theorem 3.15] tells us that $\overline{d}=d^{*}$ and $\underline{d}=d_{*}$. More generally, we see that $\overline{d}\leq d^{*}$ (hence $\underline{d}\geq d_{*}$) since any Følner sequences generates many means, but it is not clear if $d^{*}\leq\overline{d}$.
Suppose that we are given a function $f:S\rightarrow\mathbb{C}$ and $D\in\{d_{*},\underline{d},d_{\mathcal{F}}\}$ that is well-defined with respect to $S$. We write
$D-\lim_{x}f(x)=\alpha$
if for any open neighborhood $U$ of $\alpha$,
$D(\{x\in X:f(x)\in U\})=1.$
Given a topological space $Y$, a function $f:X\rightarrow Y$, and a set $A\subseteq X$, we write
$\lim_{\underset{x\in A}{x\to\infty}}f(x)=\alpha$
if for any open neighborhood $U$ of $\alpha$ the set $\{x\in A|f(x)\notin U\}$ is compact.
A group $G$ is amenable if there exists a left-invariant mean $m_{\ell}$ on $C_{r}(G)$, the space of bounded right uniformly continuous functions on $G$. If $G$ is amenable, then there exists a right-invariant mean $m_{r}$ on $C_{\ell}(G)$ given by $m_{r}(f(g))=m_{\ell}(f(g^{-1}))$, so we do not distinguish between left and right amenability for groups. If $G$ is locally compact and $\lambda$ is a Haar measure, then $G$ is amenable if and only if $(G,\lambda)$ is amenable. Furthermore, $(G,\lambda)$ is amenable if and only if $(G,\lambda)$ possesses a left-Følner net. In any topological group $G$ we have that $W(G)\subseteq C_{r}(G)\cap C_{\ell}(G)$. It follows that if $G$ is amenable, then $W(G)$ admits a left-invariant mean and a right-invariant mean, hence it admits a bi-invariant mean. However, if $G$ is not a locally compact amenable group and $m$ is a left-invariant mean on $C_{r}(G)$, then $m$ cannot in general be extended to a left-invariant mean on $C_{b}(G)$, let alone the space of bounded Borel measurable functions on $G$. For this reasons, Theorems 4.2 and 5.2 do not apply to general amenable groups.
A group $G$ is extremely amenable if every continuous action of $G$ on a compact Hausdorff space has a fixed point. If $G$ is extremely amenable, then there exists a left invariant mean $m$ on $C_{r}(G)$ that is multiplicative in the sense that $m(fg)=m(f)m(g)$ for all $f,g\in C_{r}(G)$, which justifies the name of this class of groups. The only locally compact group that is extremely amenable is the trivial group. In particular, the class of extremely amenable groups is a naturally occurring class of amenable groups that cannot be studied with Følner nets. For further discussion on extremely amenable groups the reader is referred to [38].
2.4 Representations of semigroups
Let $E$ be a Banach space and $S$ a semigroup. A representation $\pi$ of $S$ on $E$ is a weakly continuous [^5] We warn the reader that many other texts require that $\pi$ be strongly continuous. Nonetheless, for our purposes weak continuity is enough. map $\pi:S\rightarrow\mathcal{L}(E)$. If $S$ is a monoid with identity $e$, then we will assume that $\pi_{e}$ is the identity map. We note that if $\mathcal{H}$ is a Hilbert space and $U$ is a unitary representation of $S$ on $\mathcal{H}$, then $U$ is strongly continuous. The representation $\pi$ is uniformly bounded if $\sup_{s\in S}\|\pi_{s}\|<\infty$, and it is relatively weakly compact if $\{\pi_{s}\ |\ s\in S\}$ is relatively compact in the weak topology of $\mathcal{L}(E)$. A fact that we will frequently use is that if $\pi$ is a relatively weakly compact representation of $S$ on $E$, then for any $\xi\in E$ and any $f^{\prime}\in E^{\prime}$, the function $\phi(s)=f^{\prime}(\pi_{s}\xi)$ belongs to $W(S)$ (cf. [17, Theorem 2.7]). While every relatively weakly compact set of operators in $\mathcal{L}(E)$ is uniformly bounded, the converse need not be true.
The set $\mathcal{I}(E)$ of invertible bounded linear operators on $E$ is a topological group when given the strong operator topology. Recently, Banakh [3] showed that a Haar-measurable homomorphism [^6] Given two topological groups $X$ and $Y$, a map $h:X\rightarrow Y$ is Haar-measurable if for every open set $U\subseteq Y$, the set $h^{-1}(U)$ is in the completion of the Borel $\sigma$-algebra of $X$ with respect to a Haar measure of $X$. from a locally compact group to a topological group is in fact a continuous homomorphism. In particular, if $G$ is a locally compact group and $\pi:G\rightarrow\mathcal{I}(E)$ is a Haar-measurable homomorphism, then $\pi$ is strongly (hence weakly) continuous.
A matrix coefficient $\phi$ of the representation $\pi$ is a function of the form $\phi(s)=f^{\prime}(\pi_{s}\xi)$ for some $\xi\in E$ and some $f^{\prime}\in E^{\prime}$. If $\pi_{1}$ and $\pi_{2}$ are representations of $S$ on $E_{1}$ and $E_{2}$ respectively, then $\pi_{1}$ is weakly contained in $\pi_{2}$ if every matrix coefficient of $\pi_{1}$ can be approximated arbitrary well by linear combinations of matrix coefficients of $\pi_{2}$. To be more precise, $\pi_{1}$ is weakly contained in $\pi_{2}$ if for any $\xi\in E_{1}$, any $f^{\prime}\in E_{1}^{\prime}$, and any $\epsilon>0$, there exists $\eta_{1},\cdots,\eta_{k}\in E_{2}$ and $h^{\prime}_{1},\cdots,h^{\prime}_{k}\in E^{\prime}_{2}$ for which
$\sup_{s\in S}\left|f^{\prime}(\pi_{1,s}\xi)-\sum_{j=1}^{k}h^{\prime}_{j}(\pi_{2,s}\eta_{j})\right|<\epsilon.$ (3)
The representations $\pi_{1}$ and $\pi_{2}$ are weakly equivalent if $\pi_{1}$ is weakly contained in $\pi_{2}$ and $\pi_{2}$ is weakly contained in $\pi_{1}$. The representations $\pi_{1}$ and $\pi_{2}$ are equivalent if there exists a bounded linear invertible operator $T:E_{1}\rightarrow E_{2}$ for which $\pi_{2,s}T=T\pi_{1,s}$ for all $s\in S$. It is known that if $T:E_{1}\rightarrow E_{2}$ is a bounded linear invertible map, then $E_{1}$ is a Hilbert space if and only if $E_{2}$ is a Hilbert space, so a representation of $S$ on a Banach space $E$ that is not a Hilbert space cannot be equivalent to any representation of $S$ on a Hilbert space $\mathcal{H}$. However, as we shall see in Theorem 3.4, there are many representations of $S$ on a Banach space that are weakly equivalent to a representation on a Hilbert space.
[…]
2.5 Ultrafilters and the algebra of the Stone-Čech compactification
Definition 2.1.
Given a set $S$, a collection of subsets $\mathcal{U}\subseteq\mathcal{P}(S)$ is called filter if
  1. (i)
    $\emptyset\notin\mathcal{U}$ and $S\in\mathcal{U}$.
  2. (ii)
    If $A,B\in\mathcal{U}$ then $A\cap B\in\mathcal{U}$.
  3. (iii)
    If $A\in\mathcal{U}$ and $A\subseteq B$ then $B\in\mathcal{U}$.
The collection $\mathcal{U}$ is called an ultrafilter if it is a filter and for any $A\subseteq S$ we have the either $A\in\mathcal{U}$ or $A^{c}\in\mathcal{U}$.
We let $\beta S$ denote the space of ultrafilters on $S$, which is also known as the Stone-Čech compactification of $S$ when $S$ is endowed with the discrete topology. Given a topological space $X$, a function $f:S\rightarrow X$, and a point $x\in X$, we write $p-\lim_{s}f(s)=x$ if for any open neighborhood $U$ of $x$ we have $\{s\in S|f(s)\in U\}\in p$. A useful property of ultrafilters is that for any compact Hausdorff topological space $X$, any $p\in\beta S$, and any function $f:S\rightarrow X$, there exists a unique $x\in X$ for which $p-\lim_{s}f(s)=x$. We will make use of this fact for relatively weakly compact representations $\pi$ of a semigroup $S$ on a Banach space $E$, as we can then define for each $p\in\beta S$ the operator $\pi_{p}:=p-\lim_{s}\pi_{s}$.
We embed $S$ into $\beta S$ as a dense subset by identifying each $s\in S$ with the ultrafilter $e(s):=\{A\subseteq S|s\in A\}$. When $(S,\cdot)$ is a discrete semigroup, there is a unique extension of the operation $\cdot$ from $e(S)$ to all of $\beta S$ that makes $\beta S$ a compact right topological semigroup, i.e., each of the maps $R_{p}$ for $p\in\beta S$ are continuous. One can just as well extend the semigroup operation from $S$ to $\beta S$ so as to make each of the maps $L_{p}$ continuous instead of $R_{p}$, but in general the operation cannot be extended in such a way as to make $\beta S$ a semitopological semigroup. If $\mathscr{S}$ is a compact semitopological semigroup and $i:S\rightarrow\mathscr{S}$ is a continuous homomorphism whose image is dense in $\mathscr{S}$, then there exists a unique continuous surjective homomorphism $h:\beta S\rightarrow\mathscr{S}$ satisfying $h\circ e=i$, and it is given by $h(p)=p-\lim_{s}e(s)$ (cf. [29, Theorem 4.8]). It is known that $\beta S$ possesses a smallest two-sided ideal that we denote by $K(\beta S)$, and that $K(\beta S)$ possesses idempotents.
If $S$ is a left-amenable [^7] We mention that we work with a left-amenable semigroup instead of a right-amenable semigroup because we made $\beta S$ a right topological semigroup. semigroup satisfying the strong Følner condition, then we let $\triangle^{*}(S)=\{p\in\beta S|\forall\ A\in p,d^{*}(A)>0\}$. In [26, Theorem 2.8] it is shown that $\triangle^{*}(S)$ is a closed two-sided ideal in $\beta S$, hence it contains $\overline{K(\beta S)}$. An idempotent element of $\triangle^{*}(S)$ is called an essential idempotent. Essential idempotents play an important role in ergodic Ramsey theory (cf. [10, 35, 6, 11, 15]).
For further discussion about the algebra of $\beta S$ when $S$ is a discrete semigroup, the reader is referred to [29]. When mention that the algebra of Stone-Čech compactification of a semitopological semigroup $S$ has been investigated in [2], but the situation there is much more complicated, which is why we choose to restrict our attention to discrete semigroups when working with the Stone-Čech compactification.
3 The reversible part of the JdLG-decomposition
As in section 2.2 we denote by $\mathscr{T}$ a relatively weakly compact operator semigroup acting on a Banach space $E$ and we denote by $\mathscr{S}$ the weak closure of $\mathscr{T}$. We have already seen that $\mathscr{T}$ is JdLG-admissible if and only if $K(\mathscr{S})$ is a (compact topological) group. In this case there exists a unique minimal idempotent $Q\in\mathscr{S}$. The reversible part of $E$ is defined by setting $E_{r}:=\operatorname{ran}(Q)$. Letting $\mathscr{S}_{r}$ denote the restriction of $\mathscr{S}$ to $E_{r}$, it is known that $\mathscr{S}_{r}$ is a compact topological group that is (algebraically and topologically) isomorphic to $K(\mathscr{S})$ (cf. [21, Theorem 16.24]). We call a tuple $(e_{j})_{j=1}^{n}\in E^{n}$ a finite unitary system for $\mathscr{T}$ if the vectors $e_{1},...,e_{n}$ are linearly independent, the set $F:=\operatorname{lin}\{e_{1},...,e_{n}\}$ is $\mathscr{T}$-invariant, and the corresponding matrix representation $\chi$ of $\mathscr{T}$ defined by
$Te_{i}:=\sum_{j=1}^{n}\chi_{ij}(T)e_{j}\quad(\text{for }i\in\{1,...,n\},T\in\mathscr{T}),$
is unitary, i.e. for every $T\in\mathscr{T}$, $\chi(T)$ is an unitary $n\times n$ matrix. Moreover an unitary system $(e_{j})_{j=1}^{n}$ is called irreducible if $F=\operatorname{lin}\{e_{1},...,e_{n}\}$ does not contain any nontrivial $\mathscr{T}$-invariant subspaces. By [21], Theorem 16.31 the reversible part of a JdLG-admissable semigroup $\mathscr{T}$ is given by
[…]
Under an additional assumption we can characterize the reversible part as follows.
Theorem 3.1 (cf. [21], Proposition 16.29.).
Let $\mathscr{T}$ be a JdLG-admissable semigroup of bounded linear operators on a Banach space $E$. If
$\inf_{T\in\mathscr{T}}\|Ty\|>0$
for every $y\neq 0$, such that $\mathscr{T}y$ is relatively compact in $E$, then
$E_{r}=\{\xi\in E:\mathscr{T}\xi\text{ is relatively compact in }E\}.$
Clearly, Theorem 3.1 applies to any JdLG-admissible semigroup of isometries. Another case of interest is when we have representations of groups.
Corollary 3.2.
Let $G$ be a group, $E$ be a Banach space and $\pi$ a JdLG-admissible representation of $G$ on $E$. Then
$E_{r}=\{\xi\in E:\pi(G)\xi\text{ is relatively compact in }E\}.$
Proof.
Since $\pi$ is JdLG-admissible, it is weakly almost periodic, so there exists $C>0$ such that $\sup_{g\in G}\|\pi_{g}\|\leq C$. Hence, for any $x\in E$ we obtain
$\|x\|=\left\|\pi_{g}^{-1}\pi_{g}x\right\|\leq\|\pi_{g^{-1}}\|\|\pi_{g}x\|\leq C\|\pi_{g}x\|,$
and by Theorem 3.1 the result follows. ∎
Example 3.3.
To see that Corollary 3.2 does not hold for general semigroups, let $\pi$ be the representation of $\mathbb{N}$ on $E=\mathbb{C}$ given by $\pi(n)(z)=2^{-n}z$, and observe that that $E=E_{\text{aws}}=\{\xi\in E|\pi(\mathbb{N})\xi\text{ is relatively compact}\}$ and $E_{r}=\{0\}$.
Theorem 3.4.
Let $S$ be a semigroup, $E$ a (Banach) Hilbert space, $\pi$ a JdLG-admissible representation of $S$ on $E$, and $\pi_{r}$ the restriction of $\pi$ to $E_{r}$. The representation $\pi_{r}$ is (weakly) equivalent to a unitary representaiton $U$ of $S$ on a Hilbert space $\mathcal{H}$.
Proof.
Let us first assume that $E$ is a general Banach space. Let $(F_{i})_{i\in I}$ denote the collection of finite dimensional subspaces of $E_{r}$ that have an irreducible finite unitary systems $\{e_{i,j}\}_{j=1}^{n_{i}}$ as a basis. We see that for $i\neq j\in I$, $F_{i}\cap F_{j}$ is a $\pi(S)$-invariant subspace of $F_{i}$, hence $F_{i}\cap F_{j}=\{0\}$. For $i\in I$ and $1\leq j\leq n_{i}$ let $f_{i,j}^{\prime}\in E_{r}^{\prime}$ be the functional that is 0 on the complement of $F_{i}$ and on $F_{i}$ satisfies
$f_{i,j}^{\prime}\left(\sum_{k=1}^{n_{i}}c_{i,k}e_{i,k}\right)=c_{i,j}.$ (4)
Since $E_{r}$ is the smallest closed subspace containing all of the $F_{i}$, we see that the closed linear hull of $\{\{e_{i,j}\}_{j=1}^{n_{i}}\}_{i\in I}$ is $E_{r}$, and that the closed linear hull of $\{\{f_{i,j}^{\prime}\}_{j=1}^{n_{i}}\}_{i\in I}$ is $E_{r}^{\prime}$. Let $\mathcal{H}_{i}=\mathbb{C}^{n_{i}}$ for $i\in I$. For $s\in S$ and $i\in I$, let $\pi_{i,s}$ denote the $n_{i}\times n_{i}$ unitary matrix that represents the action of $\pi_{s}$ on $F_{i}$ with respect to the basis $\{e_{i,j}\}_{j=1}^{n_{i}}$. Let $U_{i,s}$ denote the same matrix as $\pi_{i,s}$, but acting on $\mathbb{C}^{n_{i}}$ with the standard basis instead of $F_{i}$. It follows that $U=\oplus_{i\in I}U_{i}$ is a unitary representation of $S$ acting on the Hilbert space $\mathcal{H}:=\oplus_{i\in I}\mathbb{C}^{n_{i}}$. We let $\phi:E_{r}\rightarrow\mathcal{H}$ be the map that identifies $\{\{e_{i,j}\}_{j=1}^{n_{i}}\}_{i\in I}$ with the canonical orthonormal basis of $\mathcal{H}$, and we observe that $\phi$ is not necessarily a bounded linear invertible map. For $\xi\in\mathcal{H}$, let $\xi^{\prime}\in\mathcal{H}^{\prime}$ be the functional given by $\xi^{\prime}(\eta)=\langle\eta,\xi\rangle$. To see that the representation $U$ is weakly equivalent to $\pi_{r}$, it suffices to observe that for any $i_{1},i_{2}\in I$ and $(j_{1},j_{2})\in[1,n_{i_{1}}]\times[1,n_{i_{2}}]$, we have $\phi(e_{i_{1},j_{1}})^{\prime}(\phi(e_{i_{2},j_{2}}))=f_{i_{1},j_{1}}^{\prime}(e_{i_{2},j_{2}})$.
Now let us consider the case in which $E$ is a Hilbert space. We see that $E=\oplus_{i\in I}F_{i}$. We recall that $\mathscr{S}_{r}:=\overline{\pi_{r}(S)}$ is a compact (hence amenable) topological group. Day [16] showed that any uniformly bounded representation of an amenable group $G$ on a Hilbert space is unitarizable, i.e., is equivalent to a unitary representation of $G$. The identity map gives a uniformly bounded representation of $\mathscr{S}_{r}$ on the Hilbert space $E$, so it is unitarizable. Since $E$ decomposes into a direct sum of finite dimensional representations of $\mathscr{S}_{r}$, any equivalent representation does the same, which yields the desired result. ∎
4 The almost weakly stable part of the JdLG-decomposition
We begin with a characterization of the almost weakly stable part of the JdLG-decomposition through the unique invariant mean on the space of weakly almost periodic functions. This characterization is related to the so-called ”flight functions" discussed in [18, 30, 40].
Lemma 4.1.
Let $S$ be a semitopological semigroup with a bi-invariant mean $m$ on $W(S)$, let $E$ be a Banach space, and let $\pi$ be a relatively weakly compact representation of $S$ on $E$. For $\xi\in E$, the following are equivalent:
  1. (i)
    $\xi\in E_{\text{aws}}$.
  2. (ii)
    For all $p>0$, and all $f^{\prime}\in E^{\prime}$, we have
    $m\left(s\mapsto|f^{\prime}(\pi_{s}\xi)|^{p}\right)=0.$ (5)
  3. (iii)
    There exist $p>0$ such that for all $f^{\prime}\in E^{\prime}$, we have
    $m\left(s\mapsto|f^{\prime}(\pi_{s}\xi)|^{p}\right)=0.$ (6)
Proof.
We will first show that (i) implies (ii). Let $(s_{i})_{i\in I}\subseteq S$ be a sequence for which $\lim_{i}\pi_{s_{i}}\xi=0$ weakly, and let $p>0$ be arbitrary. Since for each $f^{\prime}\in E^{\prime}$ we have $s\mapsto|f^{\prime}(\pi_{s}\xi)|^{p}\in\text{W}(G)$ we have that $\{|f^{\prime}(\pi_{ss_{i}}\xi)|^{p}\}_{n=1}^{\infty}$ is a relatively weakly compact set, so there exists a subnet $(s_{j})_{j\in J}$ for which $|f^{\prime}(\pi_{ss_{j}}\xi)|^{p}$ converges weakly to some function $\phi(s)$. Since weak convergence implies pointwise convergence and $\lim_{j}\pi_{s_{j}}\xi=0$ weakly, we see that $\phi$ is the 0 function, hence
$m(|f^{\prime}(\pi_{s}\xi)|^{p})=\lim_{j}m(|f^{\prime}(\pi_{ss_{j}}\xi)|^{p})=m(\lim_{j}|f^{\prime}(\pi_{ss_{j}}\xi)|^{p})=m(0)=0.$ (7)
It is clear that (ii) implies (iii), so we will now show that (iii) implies (i). Let us assume that $\xi\notin E_{\text{aws}}$ and let us first consider the case of $p=2$. Since $\pi$ is a JdLG-admissible representation of $S$, can write $\xi=\xi_{1}+\xi_{2}$ with $0\neq\xi_{1}\in E_{\text{rev}}$ and $\xi_{2}\in E_{\text{aws}}$. Since $E_{\text{rev}}$ can be further decomposed into a direct sum of irreducible finite dimensional subspaces, let us write $\xi_{1}=x_{1}+x_{2}$ with $x_{1}$ belonging to a finite dimensional irreducible subspace $E_{1}$. We may further assume without loss of generality that $S$ acts on $E_{1}$ by unitary matrices. If $\text{dim}(E_{1})=k$, then for each $v\in E_{1}$ there exists $k$ functionals $f_{1},\cdots,f_{k}$ (depending on $v$) for which $\|v\|^{2}=\sum_{j=1}^{k}|f_{j}(\pi_{s}v)|^{2}$ for all $s\in S$. It follows that
$\|x_{1}\|^{2}=m\left(\|\pi_{s}x_{1}\|^{2}\right)=m\left(\sum_{j=1}^{k}|f_{j}(\pi_{s}x_{1})|^{2}\right)=\sum_{j=1}^{k}m\left(|f_{j}(\pi_{s}x_{1})|^{2}\right),$ (8)
so for some $1\leq j\leq k$ we have
$m\left(|f_{j}(\pi_{s}x_{1})|^{2}\right)\geq\frac{\|x_{1}\|^{2}}{k}>0.$ (9)
We extend $f_{j}$ to a functional $\varphi$ on all of $E$ by defining $\varphi(v)=0$ for all $v$ in the complement of $E_{1}$, and we see that
$m\left(|\varphi(\pi_{s}\xi)|^{2}\right)>0.$ (10)
To handle the case of a general $p>0$, we let $S^{w}$ denote the weakly almost periodic compactification of $S$, and we let $i,M,$ and $\mu$ be as in Section 2.1. Let $f^{\prime}\in E^{\prime}$ be arbitrary. Since
$m\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)=M\left(i\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)\right)=\int_{S^{w}}\left|f^{\prime}(\pi_{s}\xi)\right|^{p}d\mu,$ (11)
we see that $m(|f^{\prime}(\pi_{s}\xi|^{p})=0$ if and only if $i(|f^{\prime}(\pi_{s}\xi)|)^{p}=0$ for $\mu$-a.e. $s\in S^{w}$. Lastly, we observe that $i(|f^{\prime}(\pi_{s}\xi)|)^{p}=0$ for $\mu$-a.e. $s\in S^{w}$ if and only if $i(|f^{\prime}(\pi_{s}\xi)|)^{2}=0$ for $\mu$-a.e. $s\in S^{w}$. ∎
Our next result can be seen as a generalization of [21, Theorem 16.34] (see also [31] and [22, Theorem 2.5]), and also relates the almost weakly stable vectors to the vectors from $\mathcal{H}_{w}$ in the compact-weak mixing decomposition.
Theorem 4.2.
Let $(S,\lambda)$ be a bi-amenable measured semigroup. Let $E$ be a Banach space, and let $\pi$ be a relatively weakly compact (hence JdLG-admissible) representation of $S$ on $E$. Given $\xi\in E$ and a left-invariant mean $m$ on $L^{\infty}(S,\lambda)$, (i)-(iv) are equivalent.
  1. (i)
    $\xi\in E_{\text{aws}}$, i.e., $0\in c\ell_{\sigma}(\pi(S)\xi)$.
  2. (ii)
    For some/each $p>0$, $m\left(s\mapsto\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}\right)=0\ \forall\ x^{\prime}\in E^{\prime}$.
  3. (iii)
    For any weakly open neighborhood $U$ of $0\in E$, $m\left(\left\{s\in S|\pi_{s}\xi\in U\right\}\right)=1$.
  4. (iv)
    $d_{*}-\displaystyle\lim_{g}\pi_{g}\xi=0\text{ weakly}$.
Moreover, if $S$ is locally compact, $\lambda$ is a Radon measure, and $(S,\lambda)$ possesses a left-Følner net $\mathcal{F}=(F_{i})_{i\in I}$, then (i)-(vii) are equivalent.
  1. (v)
    For some/each $p>0$, $\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}d\lambda(s)=0\ \forall\ x^{\prime}\in E^{\prime}$.
  2. (vi)
    $d_{\mathcal{F}}-\displaystyle\lim_{g}\pi_{g}\xi=0\text{ weakly}$.
  3. (vii)
    $\underline{d}-\displaystyle\lim_{g}\pi_{g}\xi=0\text{ weakly}$.
If we further assume $\mathcal{F}=(F_{n})_{n=1}^{\infty}$ is a Følner sequence, then (i)-(viii) are equivalent.
  1. (viii)
    For some/each $p>0$, $\lim_{n\rightarrow\infty}\sup_{\|x^{\prime}\|\leq 1}\frac{1}{\lambda(F_{n})}\int_{F_{n}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}d\lambda(s)=0$.
If we assume even further that $E^{\prime}$ is separable and $\lim_{n\rightarrow\infty}\lambda(F_{n})=\infty$, then (i)-(ix) are equivalent.
  1. (ix)
    There exists a set $A\subseteq S$ with $d_{\mathcal{F}}(A)=1$ such that $\lim_{\underset{s\in A}{s\to\infty}}T_{s}\xi=0$ weakly.
Proof.
As discussed in Section 2.1, $m$ restricts to a bi-invariant mean on $W(S)$. We now use Lemma 4.1 to see that for some/each $p>0$ we have
$E_{\text{aws}}$ $=\left\{\xi\in E\ |\ m\left(s\mapsto\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}\right)=0\ \forall\ x^{\prime}\in E^{\prime}\right\},$
hence (i) and (ii) are equivalent.
To see that (ii) implies (iii), let $x^{\prime}\in E^{\prime}$ and $\epsilon>0$ both be arbitrary, let $U=\{\eta\in E||\langle\eta,x^{\prime}\rangle|^{p}<\epsilon$, and observe that
$0=m\left(s\mapsto|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}\right)\geq\epsilon m\left(\mathbbm{1}_{U^{c}}(\pi_{s}\xi)\right)\text{, hence }m(\mathbbm{1}_{U}(\pi_{s}\xi))=1-m\left(\mathbbm{1}_{U^{c}}(\pi_{s}\xi)\right)=1.$ (12)
To see that (iii) implies (ii), we observe that
$m\left(s\mapsto|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}\right)\leq\epsilon m\left(\mathbbm{1}_{U}(\pi_{s}\xi)\right)+m\left(\mathbbm{1}_{U^{c}}(\pi_{s}\xi)\right)=\epsilon.$ (13)
It is clear that (iv) implies (iii). To see that (iii) implies (iv), we recall that (iii) is equivalent to (i), and (i) is independent of the choice of $m$. In particular, if (iii) holds for a particular left-invariant mean, then (iii) holds for all left-invariant means, which yields (iv).
Now let us assume that $(S,\lambda)$ admits a left-Følner net $\mathcal{F}=(F_{i})_{i\in I}$. We observe that $g(s):=|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}$ is in $W(S)$. For each $\in\mathcal{M}(\mathcal{F}), mm$ restricts to the unique bi-invariant mean on $W(s)$, so the value of $m(g)$ is independent of the choice of $m$, hence
$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}d\lambda(s)=m(g)\text{ for all }m\in\mathcal{M}(\mathcal{F}).$ (14)
In particular, the limit in the left hand side of (14) exists. Recalling that if (ii) holds for a particular left-invariant mean then it holds for all left-invariant means, we see that (ii) is equivalent to (v). Similarly, we see that (iii) is equivalent to (vi).
It is clear that (vii) implies (vi). To see that (vi) implies (vii), it suffices to observe that (vi) is equivalent to (i), so if (vi) holds for a particular left-Følner sequence $\mathcal{F}$, then (vi) holds for every left-Følner sequence $\mathcal{F}$.
Now let us assume that $\mathcal{F}=(F_{n})_{n=1}^{\infty}$ is a Følner sequence. It is now clear that (viii) implies (v), so we will proceed to show that (v) implies (viii). Since $(\pi_{s})_{s\in S}$ is uniformly bounded, we suppose without loss of generality that $S$ is a monoid and that $(\pi_{s})_{s\in S}$ consists of contractions. Then the adjoint action $(\pi_{s}^{*})_{s\in S}$ leaves the ($\sigma^{*}$-compact) dual unit ball $D^{\prime}\subseteq E^{\prime}$ invariant. Let $(T_{s})_{s\in S}$ denote the Koopman action on $C(D^{\prime})$ induced by $(\pi^{*}_{s})_{s\in S}$. For $\xi$ satisfying (ii) we let $f_{\xi}\in C(D^{\prime})$, $f_{\xi}(x^{\prime}):=\left|\langle\xi,x^{\prime}\rangle\right|^{p}$, $p>0$. We observe by assumption, that
$\lim_{n\rightarrow\infty}\frac{1}{\lambda(F_{n})}\int_{F_{n}}T_{s}f_{\xi}(x^{\prime})d\lambda(s)=\lim_{n\rightarrow\infty}\frac{1}{\lambda(F_{n})}\int_{F_{n}}f_{\xi}(\pi_{s}^{*}x^{\prime})d\lambda(s)=0$ (15)
holds pointwise. Then, by the dominated convergence theorem the convergence of 15 holds weakly, i.e.,
$\lim_{n\rightarrow\infty}\frac{1}{\lambda(F_{n})}\int_{F_{n}}\langle T_{s}f_{\xi},\mu\rangle d\lambda(s)=0$ (16)
holds for any $\mu\in\left(C(D^{\prime})\right)^{\prime}$. By Eberlein’s generalized mean ergodic theorem, weak convergence implies the convergence of 16 in norm (see, e.g., Chapter 2, Theorem 1.5 of [34]). Hence
$0=\lim_{n\rightarrow\infty}\left\|\frac{1}{\lambda(F_{n})}\int_{F_{n}}T_{s}f_{\xi}d\lambda(s)\right\|_{\infty}=\lim_{n\rightarrow\infty}\sup_{\|x^{\prime}\|\leq 1}\frac{1}{\lambda(F_{n})}\int_{F_{n}}|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}d\lambda(s).$
Now let us show that (vi) is equivalent to (ix) under additional stated assumptions. Since $d_{\mathcal{F}}(C)=0$ for every compact set $C$, we see that (ix) implies (vi), so it remains to show that (vi) implies (ix). Since $E^{\prime}$ is separable, it is first countable, so let $\{U_{n}\}_{n=1}^{\infty}$ be a neighborhood basis for $0\in E^{\prime}$. For each $n\in\mathbb{N}$, let $A_{n}=\{s\in S|\pi_{s}\xi\in U_{n}\}$ and let $m_{n}$ be such that $\lambda(A_{n}\cap F_{m})>\frac{n-1}{n}\lambda(F_{m})$ for all $m\geq m_{n}$. Assuming without loss of generality that $m_{n+1}>m_{n}$, we see that it suffices to take $A=\bigcup_{n=1}^{\infty}\bigcup_{m=m_{n}+1}^{m_{n+1}}(A_{n}\cap F_{m})$. ∎
Example 4.3.
We now recall [32, Example 4.3] to give an example showing that Lemma 4.1 does not necessarily hold if $W(S)$ only admits a left-invariant mean and not a bi-invariant mean. This example will also show that Theorem 4.2 does not necessarily hold if $S$ is a left-amenable subsemigroup of an amenable group that is not right amenable.
We consider the semigroup $(S,\cdot)$ given by $S=\{(m,n)\in\mathbb{Z}^{2}\ |\ m\geq 0,n\geq 1\}$ and $(m_{1},n_{1})\cdot(m_{2},n_{2})=(2^{n_{2}}m_{1}+m_{2},n_{1}+n_{2})$. $S$ is cancellative and left-amenable, hence it satisfies the strong Følner condition and possesses a left-Følner sequence $\mathcal{F}=(F_{i})_{i=1}^{\infty}$. Since $S\cdot(0,1)\cap S\cdot(1,1)=\emptyset$, we see that $S$ is not right amenable. Furthermore, by replacing $F_{n}$ with $F_{n}\cdot(0,1)$, we may assume without loss of generality that $F_{n}\subseteq S\cdot(0,1)$ for all $n$. We observe that $S=S\cdot(0,1)\cup S\cdot(1,1)\cup\mathbb{Z}_{\geq 0}\times\{0\}$. Now consider the function $f:S\rightarrow\{0,1\}$ given by $f(s)=1$ if and only if $s\in S\cdot(0,1)$. We see that $\{R_{s}f\}_{s\in S}=\{f,0,1,\mathbbm{1}_{A}-f\}$, where $A=S\cdot(0,1)\cup S\cdot(1,1)$, hence it is relatively weakly compact and we see that $f\in W(S)$. Now let $E=W(S)$, $\xi=f$, $\pi=R$, and $x^{\prime}$ denote evaluation at $s_{0}$ for some $s_{0}\in S$. We see that $f\in E_{\text{aws}}$ because $R_{(1,1)}f=0$. However, if $m\in\mathcal{M}(\mathcal{F})$, then
$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|d\lambda(s)=m(s\mapsto\left(R_{s}f\right)(s_{0}))=m(1)=1.$ (17)
5 The JdLG decomposition via ultrafilters
Let $E=\mathcal{H}$ be a Hilbert space, $U$ a unitary operator on $\mathcal{H}$, and $\mathcal{H}=\mathcal{H}_{c}\oplus\mathcal{H}_{w}$ the compact-weak mixing decomposition. There are various results in the literature, such as [8, Theorem 2.2], [10, Theorem 2.25], [35, Theorem 4], and [12, Section 6], that (allow one to) construct the orthogonal projection $P:\mathcal{H}\rightarrow\mathcal{H}_{c}$ as an ultrafilter limit of powers of $U$. Theorem 5.2 is motivated by these previously mentioned results and attempts to generalize them to JdLG-admissible representations on a Banach space $E$. Strictly speaking, we cannot say that Theorem 5.2 generalizes all of the aforementioned results, because they also consider various classes of polynomial subsequences that we do not consider. For the next two results, we freely use the notation and facts from Subsection 2.5.
Lemma 5.1.
Let $S$ be a discrete semigroup, let $E$ be a Banach space, let $\pi$ be a JdLG-admissible representation of $S$ on $E$, and let $E=E_{r}\oplus E_{\text{aws}}$ be the JdLG decomposition.
  1. (i)
    If $p\in\beta S$ is an idempotent, then $p-\lim_{s}\pi_{s}\xi_{r}=\xi_{r}$ for all $\xi_{r}\in E_{r}$.
  2. (ii)
    If $S$ is a bi-amenable semigroup and $q\in\triangle^{*}(S)$, then $q-\lim_{s}\pi_{s}\xi_{\text{aws}}=0$ for all $\xi_{\text{aws}}\in E_{\text{aws}}$.
Proof.
We first prove (i). Let $h:\beta S\rightarrow\mathscr{S}_{r}=K(\mathscr{S})$ denote the map $h(q)=\displaystyle q-\lim_{s}\pi_{s}\xi$ with convergence taking place in the weak topology of $E_{r}$ (instead of the weak topology of $E$). Since $\pi:S\rightarrow\mathscr{S}$ is a homomorphism with dense image in a semitopological semigroup and $\pi=h\circ e$ on $S$, we see that $h$ a continuous surjective homomorphism. In particular, $h$ sends idempotents to idempotents. Since $\mathscr{S}_{r}=K(\mathscr{S})$ is a compact topological group, so it contains a unique idempotent $Q$. Furthermore, we have already seen that $Q$ acts as the identity on $E_{r}$, so the desired result follows.
We now proceed to prove (ii). Theorem 4.2 tells us that for any weakly open neighborhood $U$ of 0, we have
$d^{*}\left(\left\{s\in S\ |\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\right)=0\text{, hence }\left\{s\in S\ |\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\notin q.$ (18)
Since $q$ is an ultrafilter, we see that $\left\{s\in S\ |\ \pi_{s}\xi_{\text{aws}}\in U\right\}\in q$. Since $U$ was arbitrary, we see that $q-\lim_{s}\pi_{s}\xi_{\text{aws}}=0$. ∎
Theorem 5.2.
Let $S$ be a discrete semigroup, let $E$ be a Banach space, and let $\pi$ be a JdLG-admissible representation of $S$ on $E$.
  1. (i)
    If $p\in\overline{K(\beta S)}$ is an idempotent, then $\pi_{p}:=p-\lim_{s}\pi_{s}=Q$, where $Q$ is the projection onto $E_{r}$ whose kernel is $E_{\text{aws}}$.
  2. (ii)
    If $S$ is a bi-amenable, then for each essential idempotent $p\in\beta S$ we have $\pi_{p}=Q$.
  3. (iii)
    If $S$ is an amenable group, and $p\in\beta S$ is of the form $qq^{-1}$ for some $q\in\triangle^{*}(S)$, then $\pi_{p}=Q$.
Proof.
We begin with the proof of (i). We let $\mathscr{S}=c\ell_{\sigma}(\pi(S))$ and we recall that $\mathscr{S}$ is a compact semitopological semigroup. Let $\phi:\beta S\rightarrow\mathscr{S}$ be given by $\phi(q)=\pi_{q}=q-\lim_{s}\pi_{s}$, and observe that as in the proof of Lemma 5.1, $\phi$ is continuous surjective homomorphism. It follows that $K(\mathscr{S})=\phi(K(\beta S))$. Since $\pi$ is a JdLG-admissible representation, we see that $K(\mathscr{S})$ is a compact topological group with a unique idempotent $Q$, which is the projection onto $E_{r}$. Since $K(\mathscr{S})$ is closed, we see that $\phi\left(\overline{K(\beta S)}\right)=K(\mathscr{S})$. Since $\phi$ is a homomorphism, $\phi$ maps idempotents to idempotents, hence $\phi$ maps all idempotents of $\overline{K(\beta S)}$ to $Q$, which completes the proof of (i).
We now turn to the proof of (ii). Let $\xi\in E$ be arbitrary and let $\xi_{r}\in E_{r}$ and $\xi_{\text{aws}}\in E_{\text{aws}}$ be such that $\xi=\xi_{r}+\xi_{\text{aws}}$. Lemma 5.1 tells us that $\pi_{p}\xi_{r}=\xi_{r}$ because $p$ is idempotent, and that $\pi_{p}\xi_{\text{aws}}=0$ because $p\in\triangle^{*}(S)$, hence $\pi_{p}\xi=\xi_{r}$, i.e., $\pi_{p}=Q$.
We proceed to the proof of (iii) in which $S=G$ is an amenable group. Let $i:G\rightarrow G$ denote the inverse map, i.e., $i(g)=g^{-1}$. The unique continuous extension $\tilde{i}:\beta G\rightarrow\beta G$ of $i$ is given by $\tilde{i}(q)=\{A\subseteq G|A^{-1}\in q\}=\{A^{-1}|A\in q\}$. Since $\triangle^{*}(S)$ is a two-sided ideal, we see that $qq^{-1}\in\triangle^{*}(S)$, hence on $E_{\text{aws}}$ we have $\pi_{qq^{-1}}=0$. Let $h:\beta G\rightarrow K(\mathscr{S})$ be as in Lemma 5.1. Since $K(\mathscr{S})$ is a topological group, multiplication is jointly continuous, hence for each $q\in\beta G$ we have
$h\left(qq^{-1}\right)$ $=h(q)h\left(q^{-1}\right)=\left(q-\lim_{g}h(g)\right)\left(q^{-1}-\lim_{g}h(g)\right)=\left(q-\lim_{g}h(g)\right)\left(q-\lim_{g}h\left(g^{-1}\right)\right)$
$=q-\lim_{g}h(g)h\left(g^{-1}\right)=q-\lim_{g}Q=Q.$
It follows that $\phi\left(qq^{-1}\right)=Q$. ∎
Remark.
We recall that the map $p\mapsto\pi_{p}$ is a homomorphism. In particular, if $p,q\in\beta S$ are such that $\pi_{p}=\pi_{q}=Q$, the $\pi_{pq}=Q$. It follows that if $S$ is a discrete amenable group and $p$ is an ultrafilter of the form $p=q_{1}q_{1}^{-1}q_{2}q_{2}^{-1}$ for some $q_{1},q_{2}\in\triangle^{*}(S)$, then $\pi_{p}=Q$. This is of significance because such ultrafilters are related to the iterated difference sets considered in [12].

### v2

@@
 Abstract
-For any JdLG-admissible representation $\pi$ of a semigroup $S$, we show that the reversible part is weakly equivalent to a unitary representation on a Hilbert space that decomposes into a direct sum of finite dimensional representations, and we give an alternative characterization of the almost weakly stable part in terms of the unique invariant mean on the space of weakly almost periodic functions.
+For any JdLG-admissible representation $\pi$ of a semigroup $S$ on a Banach space $E$, we show that the reversible part is weakly equivalent to a unitary representation on a Hilbert space that decomposes into a direct sum of finite dimensional representations, and we give an alternative characterization of the almost weakly stable part in terms of the unique invariant mean on the space of weakly almost periodic functions.
 In the case that $S$ is a bi-amenable measured semigroup, we characterize the almost weakly stable part using invariant means and averages along Følner sequences.
@@
 […]
-$\mathcal{H}_{w}$ $=\left\{\xi\in\mathcal{H}\ |\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\xi\rangle|=0\right\}=\left\{\xi\in\mathcal{H}\ |\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\eta\rangle|=0\ \forall\ \eta\in\mathcal{H}\right\}.$
+$\mathcal{H}_{w}$ $=\left\{\xi\in\mathcal{H}\ :\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\xi\rangle|=0\right\}=\left\{\xi\in\mathcal{H}\ :\ \lim_{N\rightarrow\infty}\frac{1}{N}\sum_{n=1}^{N}|\langle U^{n}\xi,\eta\rangle|=0\ \forall\ \eta\in\mathcal{H}\right\}.$
 Koopman and von Neumann [31] were the first to show that a dynamical system is weakly mixing if and only if it has no nontrivial eigenfunctions.
-A version of the compact-weak mixing decomposition for locally compact groups was found by Godement [27] (see also [40, Appendix 10.D]), and a version for amenable semigroups was found by Dye [18].
-In between Godement and Dye, the compact-weak mixing decomposition was extended to various semigroups of operators on a Banach space by Jacobs, de Leeuw, and Glicksberg [30, 17], and is now known as the Jacobs-de Leeuw-Glicksberg (JdLG) decompoisition.
+A version of the compact-weak mixing decomposition for locally compact groups was found by Godement [26] (see also [38, Appendix 10.D]), and a version for amenable semigroups was found by Dye [17].
+In between Godement and Dye, the compact-weak mixing decomposition was extended to various semigroups of operators on a Banach space by Jacobs, de Leeuw, and Glicksberg [28, 16], and is now known as the Jacobs-de Leeuw-Glicksberg (JdLG) decomposition.
 Several applications of the JdLG-decomposition can be found in the literature, e.g.
-it is used in ergodic Ramsey theory [7, Theorem 2.3], the proof of the Erdös sumset conjecture[36], the Fürstenberg-Zimmer structure theorem [19], a proof of the Wiener-Wintner theorem [1, Proposition 2.2], and more generally in the study of weighted ergodic theorems [24, 4].
-Further uses include results related to the entangled ergodic theorem [23], the Perron-Frobenius theory for completely positive operators on $W^{*}$-algebras [5], and evolution equations [25, Section V.2].
-A modern treatment of the JdLG decomposition is given in [21, Chapter 16].
+it is used in ergodic Ramsey theory [7, Theorem 2.3], the first proof of the Erdös sumset conjecture [35], the Fürstenberg-Zimmer structure theorem [18], a proof of the Wiener-Wintner theorem [1, Proposition 2.2], and more generally in the study of weighted ergodic theorems [23, 4].
+Further uses include results related to the entangled ergodic theorem [22], the Perron-Frobenius theory for completely positive operators on $W^{*}$-algebras [5], and evolution equations [24, Section V.2].
+A modern treatment of the JdLG-decomposition is given in [20, Chapter 16].
 In the present paper we will have a heavy focus on representations of semigroups, so we give a statement of the JdLG decomposition as follows.
-Theorem 1.1 (cf. [13, Theorem 6.2.18] or [28, Theorem 3.8.5]).
+Theorem 1.1 (cf. [12, Theorem 6.2.18]).
 Let $S$ be a semitopological semigroup which admits a bi-invariant mean $m$ on $W(S)$, the space of weakly almost periodic functions on $S$.
@@
 For $\mathscr{S}=c\ell_{\sigma}(\pi(S))$, the weak closure of $\pi(S)$, we have $E=E_{r}\oplus E_{\text{aws}}$, where
-$E_{r}$ $=\{\xi\in E\ |\ \forall\ u\in\mathscr{S},\exists\ v\in\mathscr{S}\text{ with }vu\xi=\xi\}\text{ and}$
-$E_{\text{aws}}$ $=\{\xi\in E\ |\ u\xi=0\text{ for some }u\in\mathscr{S}\},$
+$E_{r}$ $=\{\xi\in E\ :\ \forall\ u\in\mathscr{S},\exists\ v\in\mathscr{S}\text{ with }vu\xi=\xi\}\text{ and}$ (1)
+$E_{\text{aws}}$ $=\{\xi\in E\ :\ u\xi=0\text{ for some }u\in\mathscr{S}\},$
 and both of $E_{r}$ and $E_{\text{aws}}$ are closed $\mathscr{S}$-invariant subspaces.
@@
 We now recall the list of examples of JdLG-admissible semigroups.
-Theorem 1.2 (cf. [21, Chapter 16.3]).
+Theorem 1.2 (cf. [20, Chapter 16.3]).
 Let $E$ be a Banach space and $\mathscr{T}$ a semigroup of bounded linear operators on $E$. Then $\mathscr{T}$ is JdLG-admissible in the following cases:
@@
     $E=L^{p}(X,\mu)$ for some measure space $(X,\mu)$ and $1<p<\infty$, and $\mathscr{T}$ consists of contractions.
+Remark.
+In many cases of interest, the Banach space $E$ is reflexive.
+Consequently, it is worth recalling that a set of bounded linear operators on a reflexive Banach space $E$ is relatively weakly compact if and only if it is uniformly bounded.
 We call a representation $\pi$ of a semigroup $S$ on a Banach space $E$ JdLG-admissible if $\mathscr{T}:=(\pi_{s})_{s\in S}$ is JdLG-admissible.
 It is known that a relatively weakly compact semigroup $\mathscr{T}$ of operators is JdLG-admissible if and only if there exists a bi-invariant mean $M$ on $C_{b}(\mathscr{T})$ (cf.
-[17, Theorem 4.11]).
-If $S$ is a locally compact group, an amenable group, or a bi-amenable measured semigroup, then it has a unique invariant mean on $W(G)$.
-Consequently, the following result follows immediately from Theorem 1.1.
+[16, Theorem 4.11]).
+If $S$ is a topological group, or a bi-amenable measured semigroup, then it has a unique invariant mean on $W(G)$.
+Consequently, the next result follows immediately from Theorem 1.1.
 Corollary 1.3.
-Let $S$ be either a locally compact group, an amenable group, or a bi-amenable measured semigroup.
-If is $E$ a Banach space and $\pi$ is a relatively weakly compact representation of $S$ on $E$, then $\pi$ is JdLG-admissible.
+Let $S$ be either a topological group or a bi-amenable measured semigroup.
+If $E$ is a Banach space and $\pi$ is a relatively weakly compact representation of $S$ on $E$, then $\pi$ is JdLG-admissible.
 In particular, if $\mathscr{T}$ is a semigroup of bounded linear operators on $E$ with $c\ell_{\sigma}(\mathscr{T})=c\ell_{\sigma}(\pi(S))$, then $\mathscr{T}$ is JdLG-admissible.
@@
 Alternative characterizations of $E_{r}$ are already given in Proposition 16.29 and Theorem 16.31 of [20], and we will build upon these characterizations in Section 3.
-Alternative characterizations of $E_{\text{aws}}$ in terms of flight functions were already alluded to in [17, Corollary 5.9], [18, Theorem 1], and [28, Page 88], but were not stated explicitly prior to this paper.
-In particular, the next result follows from Theorem 3.4 and Lemma 4.1.
+Alternative characterizations of $E_{\text{aws}}$ in terms of flight functions were already alluded to in [16, Corollary 5.9] and [17, Theorem 1], but were not stated explicitly prior to this paper.
+It is now known (cf.
+[12, Theorem 4.3.13]) that a semigroup $S$ admits a bi-invariant mean $m$ on $W(S)$ if and only if we have $W(S)=\mathcal{SAP}(S)\oplus W_{0}(S)$, where $\mathcal{SAP}(S)$ is the space of strongly almost periodic functions on $S$, and $W_{0}(S)$ consists of those functions $f\in W(S)$ that have 0 in the weak closure of the orbit of their right translates, and in this case we also have that $W_{0}(S)=\{f\in W(S)\ :\ m(|f|)=0\}$.
+In particular, the next result follows from Theorem 3.6 and Lemma 4.1 and can be seen as an analogue of this classical decomposition.
 Theorem 1.4.
@@
     Let $M$ be a bi-invariant mean on $C_{b}(\mathscr{T})$. We have
-    $E_{\text{aws}}=\left\{\xi\in E\ |\ \forall\ f^{\prime}\in E^{\prime},M\left(t\mapsto\left|f^{\prime}(\pi_{t}\xi)\right|\right)=0\right\}.$ (1)
-When $S=\mathbb{N}$ in Theorem 1.1, alternative characterizations of $E_{\text{aws}}$ resembling the characterizations of $\mathcal{H}_{w}$ are stated in [21, Theorem 16.34].
+    $E_{\text{aws}}=\left\{\xi\in E\ :\ \forall\ f^{\prime}\in E^{\prime},M\left(t\mapsto\left|f^{\prime}(\pi_{t}\xi)\right|\right)=0\right\}.$ (2)
+When $S=\mathbb{N}$ in Theorem 1.1, alternative characterizations of $E_{\text{aws}}$ resembling the characterizations of $\mathcal{H}_{w}$ are stated in [20, Theorem 16.34].
 In Theorem 4.2 we extend these alternative characterizations to various classes of amenable semigroups.
-For now, we only state a particularly nice corollary of Corollary 1.3, Corollary 3.2, and Theorem 4.2.
+For now, we only state a particularly nice result following from Corollary 1.3, Corollary 3.2, and Theorem 4.2.
 Theorem 1.5.
@@
 $E_{r}$ $=\{\xi\in E:\pi(G)\xi\text{ is relatively compact in }E\}\text{, and}$
-$E_{aws}$ $=\left\{\xi\in E:\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{g}\xi,x^{\prime}\right\rangle\right|d\lambda(g)=0\text{ for all }x^{\prime}\in E^{\prime}\right\}$
-Remark.
-In many cases of interest, the Banach space $E$ is reflexive.
-Consequently, it is worth recalling that a set of bounded linear operators on a reflexive Banach space $E$ is relatively weakly compact if and only if it is uniformly bounded.
+$E_{\text{aws}}$ $=\left\{\xi\in E:\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{g}\xi,x^{\prime}\right\rangle\right|d\lambda(g)=0\text{ for all }x^{\prime}\in E^{\prime}\right\}.$
 The structure of this paper is as follows.
@@
 The second author would like to thank Uta Freiberg for funding a research visit to TU Chemnitz.
-We would like to thank Yulia Kuznetsova for providing us with Example 4.3.
 We would also like to thank Rigoberto Zelada for helpful discussions about Theorem 5.2.
@@
 2.1 Topological semigroups and weak almost periodicity
+All of the facts that we collect in this section about weakly almost periodic functions come from the classical paper of de Leeuw and Glicksberg [16].
+For futher discussions on this topic the reader is referred to [12, 13].
 A semigroup $(S,\cdot)$ is a set $S$ with an associative binary operatation $\cdot:S\times S\rightarrow S$.
 If $S$ possesses an identity element $e$, then $S$ is called a monoid.
-If $S$ is also a Hausdorff topological space and the map $\cdot$ is continuous, then $S$ is a topological semigroup [^4] We warn the reader that in [17], the term topological semigroup refers to what we call a semitopological semigroup..
+If $S$ is also a Hausdorff topological space and the map $\cdot$ is continuous, then $S$ is a topological semigroup [^5] We warn the reader that in [16] and [13], the term topological semigroup refers to what we call a semitopological semigroup..
 If $S$ is a Hausdorff topological space for which $\cdot$ is separately continuous, i.e., for each $s\in S$ the maps $L_{s}(t):=s\cdot t$ and $R_{s}(t):=t\cdot s$ are continuous, then $S$ is a semitopological semigroup.
 For the rest of this paper all semigroups are assumed to be semitopological unless it is explicitly stated otherwise.
-By abuse of notation, we let $R$ act on $C_{b}(S)$ and $L^{\infty}(S,\lambda)$ for some Borel measure $\lambda$, by $R_{s}(f)(t)=f(R_{s}(t))=f(t\cdot s)$, and similarly for $L$.
-A function $f\in C_{b}(S)$ is weakly almost periodic if $\{R_{s}f\ |\ s\in S\}$ is relatively weakly compact.
+By abuse of notation, we let $R$ act on $C_{b}(S)$ and on $L^{\infty}(S,\lambda)$ for some Borel measure $\lambda$, by $R_{s}(f)(t)=f(R_{s}(t))=f(t\cdot s)$, and similarly for $L$.
+A function $f\in C_{b}(S)$ is weakly almost periodic if $\{R_{s}f\ :\ s\in S\}$ is relatively weakly compact.
 We let $W(S)$ denote the space of weakly almost periodic functions on $S$.
-De Leeuw and Glicksberg [17, Theorem 2.7] showed that if $S$ is a compact semigroup, then $W(S)=C(S)$.
+If $S$ is a compact semigroup, then $W(S)=C(S)$.
 For $X\in\{C_{b}(S),W(S),L^{\infty}(S,\lambda)\}$, a left-invariant mean on $X$ is a positive linear functional $m:X\rightarrow\mathbb{C}$ of norm $1$ that satisfies $m(L_{s}f)=m(f)$ for all $s\in S$ and all $f\in X$, and a right-invariant mean is defined similarly.
 The semigroup $S$ admits a bi-invariant mean on $X$ if there exists a left-invariant mean $m$ on $X$ that is also a right-invariant mean.
-In [17, Theorem 5.8] it is shown that if $W(S)$ admits a bi-invariant mean, then $W(S)$ admits a unique bi-invariant mean.
-It is also known that $W(S)$ admits a bi-invariant mean if and only it it admits a left-invariant mean and a (potentially different) right-invariant mean.
-Now let $S$ be a semigroup with bi-invariant mean $m$ on $W(S)$.
-If $\mathscr{S}$ is a compact semigroup and $\pi:S\rightarrow\mathscr{S}$ is a continuous homomorphism with dense image, then $\tilde{\pi}:C(\mathscr{S})\rightarrow W(S)$ is given by $(\tilde{\pi}f)(s)=f(\pi_{s})$ is well defined and $M(f):=m(\tilde{\pi}f)$ is the bi-invariant mean on $W(\mathscr{S})$ (see [17, Lemma 5.2]).
-There exists a compact semigroup $S^{w}$, the weakly almost periodic compactification of $S$, and a continuous injective homomorphism $i:S\rightarrow S^{w}$ with dense image, such that for any $f\in W(S)$ there is a unique $f^{w}\in W(S^{w})$ satisfying $f^{w}(i(s))=f(s)$ for all $s\in S$ (cf.
-[17, Theorem 5.3]).
+If $W(S)$ admits a bi-invariant mean, then it is unique.
+It is also known that $W(S)$ admits a bi-invariant mean if and only if it admits a left-invariant mean and a (potentially different) right-invariant mean.
+Now let $S$ be a semigroup with a bi-invariant mean $m$ on $W(S)$.
+If $\mathscr{S}$ is a compact semigroup and $\pi:S\rightarrow\mathscr{S}$ is a continuous homomorphism with dense image, then $\tilde{\pi}:C(\mathscr{S})\rightarrow W(S)$ is given by $(\tilde{\pi}f)(s)=f(\pi_{s})$ is well defined and $M(f):=m(\tilde{\pi}f)$ is the bi-invariant mean on $W(\mathscr{S})=C(\mathscr{S})$.
+There exists a compact semigroup $S^{w}$, the weakly almost periodic compactification of $S$, and a continuous injective homomorphism $i:S\rightarrow S^{w}$ with dense image, such that for any $f\in W(S)$ there is a unique $f^{w}\in W(S^{w})$ satisfying $f^{w}(i(s))=f(s)$ for all $s\in S$.
 By abuse of notation, we let $i:W(S)\rightarrow W(S^{w})$ denote the map $f\mapsto f^{w}$, and we observe that $i$ is an algebra isomorphism.
@@
 All groups that we consider in this paper will be topological groups.
-De Leeuw and Glicksberg [17, Theorem 5.14] investigated which topological groups $G$ admit a bi-invariant mean on $W(G)$.
-66 years later, Ryll-Nardzewski [39] proved a fixed point theorem that allows one to deduce that all locally compact groups $G$ admit a bi-invariant mean on $W(G)$.
-For further discussions on this topic the reader is referred to [28, 13, 14].
+De Leeuw and Glicksberg [16] investigated which topological groups $G$ admit a bi-invariant mean on $W(G)$, and $6$ years later, Ryll-Nardzewski [37] proved a fixed point theorem that allows one to deduce that all topological groups $G$ admit a bi-invariant mean on $W(G)$.
 2.2 JdLG-admissible semigroups
@@
 $\mathcal{L}(E)$ is naturally a semitopological semigroup when endowed with the operation of composition as well as the weak operator topology.
-In [21, Chapter 16.3], a semigroup $\mathscr{T}\subseteq\mathcal{L}(E)$ is called JdLG-admissible if it is relatively compact and $\mathscr{S}:=\operatorname{cl}_{\sigma}(\mathscr{T})$ contains a unique minimal idempotent $Q$.
+In [20, Chapter 16.3], a semigroup $\mathscr{T}\subseteq\mathcal{L}(E)$ is called JdLG-admissible if it is relatively compact and $\mathscr{S}:=\operatorname{cl}_{\sigma}(\mathscr{T})$ contains a unique minimal idempotent $Q$.
 We will now show that this definition is equivalent to our definition.
 To this end, let us temporarily refer to the definition of JdLG-admissible from Section 1 as JdLG-decomposable in order to avoid confusion.
-If $\mathscr{T}$ is a relatively weakly compact semigroup, then [21, Theorem 16.5] says that the following are equivalent:
+If $\mathscr{T}$ is a relatively weakly compact semigroup, then [20, Theorem 16.5] says that the following are equivalent:
   1. 1.
@@
     $\mathscr{S}$ has a unique minimal right ideal and a unique minimal left ideal.
-[21, Theorem 16.3] tells us that every minimal right ideal of $\mathscr{S}$ is closed, and [21, Lemma 16.4] tells us that $K(\mathscr{S})$ is the union of all minimal right ideals of $\mathscr{S}$.
+[20, Theorem 16.3] tells us that every minimal right ideal of $\mathscr{S}$ is closed, and [20, Lemma 16.4] tells us that $K(\mathscr{S})$ is the union of all minimal right ideals of $\mathscr{S}$.
 It follows that when $\mathscr{T}$ is JdLG-admissible, $K(\mathscr{S})$ is a closed (hence compact) subsemigroup of $\mathscr{S}$.
-[17, Theorem 2.1] says that a compact semitopological semigroup that is algebraically a group must be a topological group.
+[16, Theorem 2.1] says that a compact semitopological semigroup that is algebraically a group must be a topological group.
 It follows that $\mathscr{T}$ is JdLG-admissible if and only if $K(\mathscr{S})$ is a compact topological group.
-[17, Theorem 4.11] states that $K(\mathscr{S})$ is a compact topological group if and only if $\mathscr{T}$ is JdLG-decomposable.
+[16, Theorem 4.11] states that $K(\mathscr{S})$ is a compact topological group if and only if $\mathscr{T}$ is JdLG-decomposable.
 Putting everything together, we see that $\mathscr{T}$ is JdLG-decomposable if and only if it is JdLG-admissible.
@@
 $\lim_{i}\frac{\lambda(sF_{i}\triangle F_{i})}{\lambda(F_{i})}=0\text{ for all }s\in S.$ (3)
-$\mathcal{F}$ is a two-sided Følner net if it is sumltaneously a left-Følner net and a right-Følner net.
+$\mathcal{F}$ is a two-sided Følner net if it is simultaneously a left-Følner net and a right-Følner net.
 If $S$ possesses a left/right/two-sided Følner net $\mathcal{F}=(F_{i})_{i\in I}$ then $S$ is left/right/bi amenable.
@@
 It is also well known that if $S$ is a discrete abelian semigroup, then $S$ possesses a two-sided Følner net.
-We refer the reader to [37, Chapter 4.22] for further discussion on Følner conditions and amenability for discrete semigroups.
+We refer the reader to [36, Chapter 4.22] for further discussion on Følner conditions and amenability for discrete semigroups.
 Now suppose that $S$ is a left-amenable semigroup. We define the upper (left) and lower (left) Banach densities on $S$ by
@@
 $\underline{d}(A)$ $=\inf\left\{\overline{d}_{\mathcal{F}}(A)|\mathcal{F}\text{ is a (left) F\o lner net in }S\right\}=1-\overline{d}(A^{c}).$
-If $S$ is a discrete semigroup and $\lambda$ is the counting measure, then [26, Theorem 3.15] tells us that $\overline{d}=d^{*}$ and $\underline{d}=d_{*}$.
+If $S$ is a discrete semigroup and $\lambda$ is the counting measure, then [25, Theorem 3.15] tells us that $\overline{d}=d^{*}$ and $\underline{d}=d_{*}$.
 More generally, we see that $\overline{d}\leq d^{*}$ (hence $\underline{d}\geq d_{*}$) since any Følner sequences generates many means, but it is not clear if $d^{*}\leq\overline{d}$.
 Suppose that we are given a function $f:S\rightarrow\mathbb{C}$ and $D\in\{d_{*},\underline{d},d_{\mathcal{F}}\}$ that is well-defined with respect to $S$. We write
-$D-\lim_{x}f(x)=\alpha$
+$D-\lim_{s}f(s)=\alpha$
 if for any open neighborhood $U$ of $\alpha$,
-$D(\{x\in X:f(x)\in U\})=1.$
-Given a topological space $Y$, a function $f:X\rightarrow Y$, and a set $A\subseteq X$, we write
-$\lim_{\underset{x\in A}{x\to\infty}}f(x)=\alpha$
-if for any open neighborhood $U$ of $\alpha$ the set $\{x\in A|f(x)\notin U\}$ is compact.
-A group $G$ is amenable if there exists a left-invariant mean $m_{\ell}$ on $C_{r}(G)$, the space of bounded right uniformly continuous functions on $G$.
-If $G$ is amenable, then there exists a right-invariant mean $m_{r}$ on $C_{\ell}(G)$ given by $m_{r}(f(g))=m_{\ell}(f(g^{-1}))$, so we do not distinguish between left and right amenability for groups.
-If $G$ is locally compact and $\lambda$ is a Haar measure, then $G$ is amenable if and only if $(G,\lambda)$ is amenable.
-Furthermore, $(G,\lambda)$ is amenable if and only if $(G,\lambda)$ possesses a left-Følner net.
-In any topological group $G$ we have that $W(G)\subseteq C_{r}(G)\cap C_{\ell}(G)$.
-It follows that if $G$ is amenable, then $W(G)$ admits a left-invariant mean and a right-invariant mean, hence it admits a bi-invariant mean.
-However, if $G$ is not a locally compact amenable group and $m$ is a left-invariant mean on $C_{r}(G)$, then $m$ cannot in general be extended to a left-invariant mean on $C_{b}(G)$, let alone the space of bounded Borel measurable functions on $G$.
-For this reasons, Theorems 4.2 and 5.2 do not apply to general amenable groups.
-A group $G$ is extremely amenable if every continuous action of $G$ on a compact Hausdorff space has a fixed point.
-If $G$ is extremely amenable, then there exists a left invariant mean $m$ on $C_{r}(G)$ that is multiplicative in the sense that $m(fg)=m(f)m(g)$ for all $f,g\in C_{r}(G)$, which justifies the name of this class of groups.
-The only locally compact group that is extremely amenable is the trivial group.
-In particular, the class of extremely amenable groups is a naturally occurring class of amenable groups that cannot be studied with Følner nets.
-For further discussion on extremely amenable groups the reader is referred to [38].
+$D(\{s\in S\ :\ f(s)\in U\})=1.$
+Given a topological space $Y$, a function $f:S\rightarrow Y$, and a set $A\subseteq S$, we write
+$\lim_{\underset{s\in A}{s\to\infty}}f(s)=\alpha$
+if for any open neighborhood $U$ of $\alpha$ the set $\{s\in A\ :\ f(s)\notin U\}$ is contained in a compact set.
 2.4 Representations of semigroups
-Let $E$ be a Banach space and $S$ a semigroup.
+Let $E$ be a Banach space, $\mathcal{L}(E)$ the space of bounded linear operators on $E$, and $S$ a semigroup.
 A representation $\pi$ of $S$ on $E$ is a weakly continuous [^6] We warn the reader that many other texts require that $\pi$ be strongly continuous.
@@
 We note that if $\mathcal{H}$ is a Hilbert space and $U$ is a unitary representation of $S$ on $\mathcal{H}$, then $U$ is strongly continuous.
-The representation $\pi$ is uniformly bounded if $\sup_{s\in S}\|\pi_{s}\|<\infty$, and it is relatively weakly compact if $\{\pi_{s}\ |\ s\in S\}$ is relatively compact in the weak topology of $\mathcal{L}(E)$.
+The representation $\pi$ is uniformly bounded if $\sup_{s\in S}\|\pi_{s}\|<\infty$, and it is relatively weakly compact if $\{\pi_{s}\ :\ s\in S\}$ is relatively compact in the weak topology of $\mathcal{L}(E)$.
 A fact that we will frequently use is that if $\pi$ is a relatively weakly compact representation of $S$ on $E$, then for any $\xi\in E$ and any $f^{\prime}\in E^{\prime}$, the function $\phi(s)=f^{\prime}(\pi_{s}\xi)$ belongs to $W(S)$ (cf.
-[17, Theorem 2.7]).
+[16, Theorem 2.7]).
+Since $W(S)$ is a $C^{*}$-algebra, it follows that $|\phi|^{p}\in W(S)$ as well for every $p>0$.
 While every relatively weakly compact set of operators in $\mathcal{L}(E)$ is uniformly bounded, the converse need not be true.
@@
 If $\pi_{1}$ and $\pi_{2}$ are representations of $S$ on $E_{1}$ and $E_{2}$ respectively, then $\pi_{1}$ is weakly contained in $\pi_{2}$ if every matrix coefficient of $\pi_{1}$ can be approximated arbitrary well by linear combinations of matrix coefficients of $\pi_{2}$.
-To be more precise, $\pi_{1}$ is weakly contained in $\pi_{2}$ if for any $\xi\in E_{1}$, any $f^{\prime}\in E_{1}^{\prime}$, and any $\epsilon>0$, there exists $\eta_{1},\cdots,\eta_{k}\in E_{2}$ and $h^{\prime}_{1},\cdots,h^{\prime}_{k}\in E^{\prime}_{2}$ for which
+To be more precise, $\pi_{1}$ is weakly contained in $\pi_{2}$ if for any $\xi\in E_{1}$, any $f^{\prime}\in E_{1}^{\prime}$, and any $\epsilon>0$, there exist $\eta_{1},\cdots,\eta_{k}\in E_{2}$ and $h^{\prime}_{1},\cdots,h^{\prime}_{k}\in E^{\prime}_{2}$ for which
 $\sup_{s\in S}\left|f^{\prime}(\pi_{1,s}\xi)-\sum_{j=1}^{k}h^{\prime}_{j}(\pi_{2,s}\eta_{j})\right|<\epsilon.$ (4)
@@
 The representations $\pi_{1}$ and $\pi_{2}$ are equivalent if there exists a bounded linear invertible operator $T:E_{1}\rightarrow E_{2}$ for which $\pi_{2,s}T=T\pi_{1,s}$ for all $s\in S$.
-It is known that if $T:E_{1}\rightarrow E_{2}$ is a bounded linear invertible map, then $E_{1}$ is a Hilbert space if and only if $E_{2}$ is a Hilbert space, so a representation of $S$ on a Banach space $E$ that is not a Hilbert space cannot be equivalent to any representation of $S$ on a Hilbert space $\mathcal{H}$.
-However, as we shall see in Theorem 3.4, there are many representations of $S$ on a Banach space that are weakly equivalent to a representation on a Hilbert space.
 […]
 2.5 Ultrafilters and the algebra of the Stone-Čech compactification
+Is $S$ is a set, we denote by $\mathcal{P}(S)$ its power set.
 Definition 2.1.
@@
 If $\mathscr{S}$ is a compact semitopological semigroup and $i:S\rightarrow\mathscr{S}$ is a continuous homomorphism whose image is dense in $\mathscr{S}$, then there exists a unique continuous surjective homomorphism $h:\beta S\rightarrow\mathscr{S}$ satisfying $h\circ e=i$, and it is given by $h(p)=p-\lim_{s}e(s)$ (cf.
-[29, Theorem 4.8]).
-It is known that $\beta S$ possesses a smallest two-sided ideal that we denote by $K(\beta S)$, and that $K(\beta S)$ possesses idempotents.
+[27, Theorem 4.8]).
+It is known that $\beta S$ possesses the smallest two-sided ideal that we denote by $K(\beta S)$, and that $K(\beta S)$ possesses idempotents.
 If $S$ is a left-amenable [^8] We mention that we work with a left-amenable semigroup instead of a right-amenable semigroup because we made $\beta S$ a right topological semigroup.
 semigroup satisfying the strong Følner condition, then we let $\triangle^{*}(S)=\{p\in\beta S|\forall\ A\in p,d^{*}(A)>0\}$.
-In [26, Theorem 2.8] it is shown that $\triangle^{*}(S)$ is a closed two-sided ideal in $\beta S$, hence it contains $\overline{K(\beta S)}$.
+In [25, Theorem 2.8] it is shown that $\triangle^{*}(S)$ is a closed two-sided ideal in $\beta S$, hence it contains $\overline{K(\beta S)}$.
 An idempotent element of $\triangle^{*}(S)$ is called an essential idempotent.
@@
 For further discussion about the algebra of $\beta S$ when $S$ is a discrete semigroup, the reader is referred to [27].
-When mention that the algebra of Stone-Čech compactification of a semitopological semigroup $S$ has been investigated in [2], but the situation there is much more complicated, which is why we choose to restrict our attention to discrete semigroups when working with the Stone-Čech compactification.
+We mention that the algebra of Stone-Čech compactification of a semitopological semigroup $S$ has been investigated in [2], but the situation there is much more complicated, which is why we choose to restrict our attention to discrete semigroups when working with the Stone-Čech compactification.
 3 The reversible part of the JdLG-decomposition
-As in section 2.2 we denote by $\mathscr{T}$ a relatively weakly compact operator semigroup acting on a Banach space $E$ and we denote by $\mathscr{S}$ the weak closure of $\mathscr{T}$.
+As in section 2.2, we denote by $\mathscr{T}$ a relatively weakly compact operator semigroup acting on a Banach space $E$, and we denote by $\mathscr{S}$ the weak closure of $\mathscr{T}$.
 We have already seen that $\mathscr{T}$ is JdLG-admissible if and only if $K(\mathscr{S})$ is a (compact topological) group.
 In this case there exists a unique minimal idempotent $Q\in\mathscr{S}$.
-The reversible part of $E$ is defined by setting $E_{r}:=\operatorname{ran}(Q)$.
+The reversible part of $E$ is defined by setting $E_{r}:=\operatorname{ran}(Q)$, and as discussed above, it is also described by (1).
 Letting $\mathscr{S}_{r}$ denote the restriction of $\mathscr{S}$ to $E_{r}$, it is known that $\mathscr{S}_{r}$ is a compact topological group that is (algebraically and topologically) isomorphic to $K(\mathscr{S})$ (cf.
-[21, Theorem 16.24]).
+[20, Theorem 16.24]).
 We call a tuple $(e_{j})_{j=1}^{n}\in E^{n}$ a finite unitary system for $\mathscr{T}$ if the vectors $e_{1},...,e_{n}$ are linearly independent, the set $F:=\operatorname{lin}\{e_{1},...,e_{n}\}$ is $\mathscr{T}$-invariant, and the corresponding matrix representation $\chi$ of $\mathscr{T}$ defined by
@@
 for every $T\in\mathscr{T}$, $\chi(T)$ is an unitary $n\times n$ matrix.
-Moreover an unitary system $(e_{j})_{j=1}^{n}$ is called irreducible if $F=\operatorname{lin}\{e_{1},...,e_{n}\}$ does not contain any nontrivial $\mathscr{T}$-invariant subspaces.
+If $(e_{j})_{j=1}^{n}$ is a finite unitary system for $\mathscr{T}$ with the corresponding matrix representation $\chi$, then for each $1\leq i,j\leq n$ the map $T\mapsto\chi_{ij}(T)$ is continuous.
+A unitary system $(e_{j})_{j=1}^{n}$ is called irreducible if $F=\operatorname{lin}\{e_{1},...,e_{n}\}$ does not contain any nontrivial $\mathscr{T}$-invariant subspaces.
 By [20], Theorem 16.31 the reversible part of a JdLG-admissable semigroup $\mathscr{T}$ is given by
 […]
-Under an additional assumption we can characterize the reversible part as follows.
+Under an additional assumption one can characterize the reversible part as follows.
 Theorem 3.1 (cf. [20], Proposition 16.29.).
@@
 Proof.
-Since $\pi$ is JdLG-admissible, it is weakly almost periodic, so there exists $C>0$ such that $\sup_{g\in G}\|\pi_{g}\|\leq C$. Hence, for any $x\in E$ we obtain
+Since $\pi$ is JdLG-admissible, it is weakly relatively compact, thus uniformly bounded, so there exists $C>0$ such that $\sup_{g\in G}\|\pi_{g}\|\leq C$. Hence, for any $x\in E$ we obtain
 $\|x\|=\left\|\pi_{g}^{-1}\pi_{g}x\right\|\leq\|\pi_{g^{-1}}\|\|\pi_{g}x\|\leq C\|\pi_{g}x\|,$
@@
 To see that Corollary 3.2 does not hold for general semigroups, let $\pi$ be the representation of $\mathbb{N}$ on $E=\mathbb{C}$ given by $\pi(n)(z)=2^{-n}z$, and observe that that $E=E_{\text{aws}}=\{\xi\in E|\pi(\mathbb{N})\xi\text{ is relatively compact}\}$ and $E_{r}=\{0\}$.
+The following proposition presents another particular case of JdLG-admissible semigroups.
+Proposition 3.4.
+Let $E$ be a Banach space and let $\mathscr{S}$ be a weakly compact semigroup of bounded linear operators on $E$ such that every $\mathscr{S}$-orbit contains a unique fixed point.
+Then $\mathscr{S}$ is JdLG-admissible.
+Moreover, $E_{r}$ is the set of fixed vectors of $\mathscr{S}$.
+Proof.
+Let $E_{1}$ denote the set of fixed points of $\mathscr{S}$, which is a closed linear subspace of $E$.
+For any $\xi\in E$ let $\xi_{1}=u_{1}\xi$ be the fixed point in $\mathscr{S}\xi$.
+Then $u_{1}(\xi-\xi_{1})=0$, so $\xi_{0}:=\xi-\xi_{1}\in E_{\text{aws}}$, which shows that $E=E_{1}+E_{\text{aws}}$.
+The set $E_{\text{aws}}$ is also a linear subspace: if $\xi_{1},\xi_{2}\in E_{\text{aws}}$ and $u_{1}\xi_{1}=0$, then $u_{1}(\xi_{1}+\xi_{2})=u_{1}\xi_{2}$; by assumption, $\mathscr{S}u_{1}\xi_{2}$ contains a fixed point $vu_{1}\xi_{2}$, but it will also be a fixed point in $\mathscr{S}\xi_{2}$ which is 0, so that $vu_{1}(\xi_{1}+\xi_{2})=0$ and $\xi_{1}+\xi_{2}\in E_{\text{aws}}$.
+(Stability under multiplication by scalars is obvious.) The intersection of these subspaces is clearly 0, so we have a direct sum decomposition.
+The space $E_{1}$ is in fact $E_{r}$ as defined in Theorem 1.1: if $\xi\in E_{r}$ and $\xi=\xi_{1}+\xi_{0}$ is its decomposition as above, then choose $u\in\mathscr{S}$ such that $u\xi_{0}=0$.
+For any $v\in\mathscr{S}$ we have then $vu\xi=\xi_{1}$, so choosing $v$ such that $vu\xi=\xi$ we obtain $\xi=\xi_{1}$ and thus $E_{r}\subset E_{1}$.
+The converse inclusion is obvious: (1) holds with any $u,v\in\mathscr{S}$.
+∎
+Example 3.5.
+If the fixed point is not unique, the statement may not hold.
+Consider the semigroup of two elements $S=\{a,b\}$ with multiplication $a^{2}=ba=a$, $b^{2}=ab=b$.
+It acts on $E=C(S)$ by right translations, so that $R_{a}f\equiv f(a)$ and $R_{b}f\equiv f(b)$ for every $f$.
+A function is fixed under this action if and only if it is constant, and is in the set $E_{\text{aws}}$ if it vanishes at $a$ or at $b$.
+But this set is not a linear subspace, and the decomposition in $E_{r}+E_{\text{aws}}$, though possible, is not unique.
+Proposition 3.4 applies to the following special semigroups not covered, in general, by Theorem 1.1: $\mathscr{S}$ is the weak closure of $\pi(P)$ where $P$ is the semigroup of normal states of a quantum semigroup $M$, and $\pi$ its so called $\ell_{2}$-bounded representation on a reflexive Banach space (see the proof of Proposition 3.1 in [33]).
+One example of such a semigroup $P$ is the set of probability measures on a locally compact group $G$, with convolution as the semigroup operation.
+This particular semigroup is known to have a two-sided invariant mean on $W(S)$ and thus falls into assumptions of Theorem 1.1, but in general this is not the case.
 Theorem 3.6.
 Let $S$ be a semigroup, $E$ a (Banach) Hilbert space, $\pi$ a JdLG-admissible representation of $S$ on $E$, and $\pi_{r}$ the restriction of $\pi$ to $E_{r}$.
-The representation $\pi_{r}$ is (weakly) equivalent to a unitary representaiton $U$ of $S$ on a Hilbert space $\mathcal{H}$.
+The representation $\pi_{r}$ is (weakly) equivalent to a unitary representation $U$ of $S$ on a Hilbert space $\mathcal{H}$.
 Proof.
-Let us first assume that $E$ is a general Banach space.
-Let $(F_{i})_{i\in I}$ denote the collection of finite dimensional subspaces of $E_{r}$ that have an irreducible finite unitary systems $\{e_{i,j}\}_{j=1}^{n_{i}}$ as a basis.
-We see that for $i\neq j\in I$, $F_{i}\cap F_{j}$ is a $\pi(S)$-invariant subspace of $F_{i}$, hence $F_{i}\cap F_{j}=\{0\}$.
-For $i\in I$ and $1\leq j\leq n_{i}$ let $f_{i,j}^{\prime}\in E_{r}^{\prime}$ be the functional that is 0 on the complement of $F_{i}$ and on $F_{i}$ satisfies
-$f_{i,j}^{\prime}\left(\sum_{k=1}^{n_{i}}c_{i,k}e_{i,k}\right)=c_{i,j}.$ (4)
-Since $E_{r}$ is the smallest closed subspace containing all of the $F_{i}$, we see that the closed linear hull of $\{\{e_{i,j}\}_{j=1}^{n_{i}}\}_{i\in I}$ is $E_{r}$, and that the closed linear hull of $\{\{f_{i,j}^{\prime}\}_{j=1}^{n_{i}}\}_{i\in I}$ is $E_{r}^{\prime}$.
-Let $\mathcal{H}_{i}=\mathbb{C}^{n_{i}}$ for $i\in I$.
-For $s\in S$ and $i\in I$, let $\pi_{i,s}$ denote the $n_{i}\times n_{i}$ unitary matrix that represents the action of $\pi_{s}$ on $F_{i}$ with respect to the basis $\{e_{i,j}\}_{j=1}^{n_{i}}$.
-Let $U_{i,s}$ denote the same matrix as $\pi_{i,s}$, but acting on $\mathbb{C}^{n_{i}}$ with the standard basis instead of $F_{i}$.
-It follows that $U=\oplus_{i\in I}U_{i}$ is a unitary representation of $S$ acting on the Hilbert space $\mathcal{H}:=\oplus_{i\in I}\mathbb{C}^{n_{i}}$.
-We let $\phi:E_{r}\rightarrow\mathcal{H}$ be the map that identifies $\{\{e_{i,j}\}_{j=1}^{n_{i}}\}_{i\in I}$ with the canonical orthonormal basis of $\mathcal{H}$, and we observe that $\phi$ is not necessarily a bounded linear invertible map.
-For $\xi\in\mathcal{H}$, let $\xi^{\prime}\in\mathcal{H}^{\prime}$ be the functional given by $\xi^{\prime}(\eta)=\langle\eta,\xi\rangle$.
-To see that the representation $U$ is weakly equivalent to $\pi_{r}$, it suffices to observe that for any $i_{1},i_{2}\in I$ and $(j_{1},j_{2})\in[1,n_{i_{1}}]\times[1,n_{i_{2}}]$, we have $\phi(e_{i_{1},j_{1}})^{\prime}(\phi(e_{i_{2},j_{2}}))=f_{i_{1},j_{1}}^{\prime}(e_{i_{2},j_{2}})$.
+Let us first consider the case in which $E$ is a Banach space.
+Let $X_{n}$ denote the set of matrix representations $\chi$ of $\pi(S)$ corresponding to some unitary system $(e_{j})_{j=1}^{n}$ in $E_{r}$, and let $X=\bigcup_{n=1}^{\infty}X_{n}$.
+For each $n\in\mathbb{N}$ and $\chi\in X_{n}$, let $\mathcal{H}_{\chi}=\mathbb{C}^{n}$, and observe that $\tilde{\chi}:=\chi\circ\pi$ can be viewed as a continuous representation of $S$ on $\mathcal{H}_{\chi}$.
+Let $\mathcal{H}=\displaystyle\oplus_{\chi\in X}\mathcal{H}_{\chi}$ be the Hilbert space direct sum of these spaces, and observe that $U=\oplus_{\chi\in X}\tilde{\chi}$ is a unitary representation of $S$ on $\mathcal{H}$ that decomposes into a direct sum of finite dimensional representations.
+We now show that $U$ is weakly contained in $\pi_{r}$.
+Every coefficient $\phi(s)=f^{\prime}(U_{s}\xi)$ of $U$ is approximated (uniformly on $S$) by coefficients with $\xi$ supported only in a finite number of $\mathcal{H}_{\chi}$, and by linearity, it is sufficient to consider only $\phi$ such that $\xi\in\mathcal{H}_{\chi}$ with a single $\chi\in X$.
+Fix $\chi\in X$, and let $(e_{j})_{j=1}^{n}$ be a corresponding unitary system for $\pi_{r}(S)$.
+For $F=\text{lin}\{e_{1},\cdots,e_{n}\}$ and for $1\leq i\leq n$ we define $f_{i}^{\prime}\in F^{\prime}$ by
+$f_{i}^{\prime}\left(\sum_{j=1}^{n}c_{j}e_{j}\right)=c_{i},$ (6)
+and extend $f_{i}^{\prime}$ to all of $E_{r}$ by the Hahn-Banach Theorem. Then we see that $\tilde{\chi}_{ij}(s)=f_{j}^{\prime}(\pi(s)e_{i})$, which yields the desired result.
+We will now show that $\pi_{r}$ is weakly contained in $U$. Let $(e_{j})_{j=1}^{n}$ be a finite unitary system in $E_{r}$ and let $f^{\prime}\in E^{\prime}$. We see that
+$f^{\prime}\left(\pi(s)e_{i}\right)=\sum_{j=1}^{n}\chi_{ij}(\pi(s))f^{\prime}(e_{j}),$ (7)
+so the matrix coefficient $s\mapsto f^{\prime}(\pi(s)e_{i})$ is a linear combination of matrix coefficients of $U$. Equality (5) now yields the desired result.
 Now let us consider the case in which $E$ is a Hilbert space.
-We see that $E=\oplus_{i\in I}F_{i}$.
-We recall that $\mathscr{S}_{r}:=\overline{\pi_{r}(S)}$ is a compact (hence amenable) topological group.
+Let $\mathscr{S}$ be the closure of $\pi(S)$ in the weak topology of $\mathcal{L}(E)$, let $\mathscr{S}_{r}$ denote the restriction of $\mathscr{S}$ to $E_{r}$, and observe that $\mathscr{S}_{r}$ is the weak closure of $\pi_{r}(S)$.
+We have already seen that $\mathscr{S}_{r}$ is a compact (hence amenable) topological group.
 Day [15] showed that any uniformly bounded representation of an amenable group $G$ on a Hilbert space is unitarizable, i.e., is equivalent to a unitary representation of $G$.
-The identity map gives a uniformly bounded representation of $\mathscr{S}_{r}$ on the Hilbert space $E$, so it is unitarizable.
-Since $E$ decomposes into a direct sum of finite dimensional representations of $\mathscr{S}_{r}$, any equivalent representation does the same, which yields the desired result.
+Applying this to $G=\mathscr{S}_{r}$, we get that the identity map is equivalent to a unitary representation $U^{\prime}$ of $\mathscr{S}_{r}$ on a Hilbert space $\mathcal{H}$.
+The Peter-Weyl Theorem tells us that any unitary representation of a compact group decomposes into a direct sum of finite dimensional irreducible representations.
+The desired result follows from the observation that $\pi_{r}$ is equivalent to the unitary representation $U=U^{\prime}\circ\pi_{r}$.
 ∎
@@
 We begin with a characterization of the almost weakly stable part of the JdLG-decomposition through the unique invariant mean on the space of weakly almost periodic functions.
-This characterization is related to the so-called ”flight functions" discussed in [18, 30, 40].
+This characterization is related to the flight functions discussed in [38, Appendix 10].
 Lemma 4.1.
@@
   3. (iii)
-    There exist $p>0$ such that for all $f^{\prime}\in E^{\prime}$, we have
-    $m\left(s\mapsto|f^{\prime}(\pi_{s}\xi)|^{p}\right)=0.$ (6)
+    There exist $p>0$ such that for all $f^{\prime}\in E^{\prime}$ Equation (8) holds.
 Proof.
 We will first show that (i) implies (ii).
-Let $(s_{i})_{i\in I}\subseteq S$ be a sequence for which $\lim_{i}\pi_{s_{i}}\xi=0$ weakly, and let $p>0$ be arbitrary.
-Since for each $f^{\prime}\in E^{\prime}$ we have $s\mapsto|f^{\prime}(\pi_{s}\xi)|^{p}\in\text{W}(G)$ we have that $\{|f^{\prime}(\pi_{ss_{i}}\xi)|^{p}\}_{n=1}^{\infty}$ is a relatively weakly compact set, so there exists a subnet $(s_{j})_{j\in J}$ for which $|f^{\prime}(\pi_{ss_{j}}\xi)|^{p}$ converges weakly to some function $\phi(s)$.
+Let $(s_{i})_{i\in I}\subseteq S$ be a net for which $\lim_{i}\pi_{s_{i}}\xi=0$ weakly, and let $p>0$ be arbitrary.
+Since for each $f^{\prime}\in E^{\prime}$ we have $s\mapsto|f^{\prime}(\pi_{s}\xi)|^{p}\in\text{W}(S)$ we have that $\{|f^{\prime}(\pi_{ss_{i}}\xi)|^{p}\}_{i\in I}$ is a relatively weakly compact set, so there exists a subnet $(s_{j})_{j\in J}$ for which $|f^{\prime}(\pi_{ss_{j}}\xi)|^{p}$ converges weakly to some function $\phi(s)$.
 Since weak convergence implies pointwise convergence and $\lim_{j}\pi_{s_{j}}\xi=0$ weakly, we see that $\phi$ is the 0 function, hence
@@
 It is clear that (ii) implies (iii), so we will now show that (iii) implies (i).
-Let us assume that $\xi\notin E_{\text{aws}}$ and let us first consider the case of $p=2$.
-Since $\pi$ is a JdLG-admissible representation of $S$, can write $\xi=\xi_{1}+\xi_{2}$ with $0\neq\xi_{1}\in E_{\text{rev}}$ and $\xi_{2}\in E_{\text{aws}}$.
-Since $E_{\text{rev}}$ can be further decomposed into a direct sum of irreducible finite dimensional subspaces, let us write $\xi_{1}=x_{1}+x_{2}$ with $x_{1}$ belonging to a finite dimensional irreducible subspace $E_{1}$.
-We may further assume without loss of generality that $S$ acts on $E_{1}$ by unitary matrices.
-If $\text{dim}(E_{1})=k$, then for each $v\in E_{1}$ there exists $k$ functionals $f_{1},\cdots,f_{k}$ (depending on $v$) for which $\|v\|^{2}=\sum_{j=1}^{k}|f_{j}(\pi_{s}v)|^{2}$ for all $s\in S$.
-It follows that
-$\|x_{1}\|^{2}=m\left(\|\pi_{s}x_{1}\|^{2}\right)=m\left(\sum_{j=1}^{k}|f_{j}(\pi_{s}x_{1})|^{2}\right)=\sum_{j=1}^{k}m\left(|f_{j}(\pi_{s}x_{1})|^{2}\right),$ (8)
-so for some $1\leq j\leq k$ we have
-$m\left(|f_{j}(\pi_{s}x_{1})|^{2}\right)\geq\frac{\|x_{1}\|^{2}}{k}>0.$ (9)
-We extend $f_{j}$ to a functional $\varphi$ on all of $E$ by defining $\varphi(v)=0$ for all $v$ in the complement of $E_{1}$, and we see that
-$m\left(|\varphi(\pi_{s}\xi)|^{2}\right)>0.$ (10)
-To handle the case of a general $p>0$, we let $S^{w}$ denote the weakly almost periodic compactification of $S$, and we let $i,M,$ and $\mu$ be as in Section 2.1.
-Let $f^{\prime}\in E^{\prime}$ be arbitrary.
-Since
-$m\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)=M\left(i\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)\right)=\int_{S^{w}}\left|f^{\prime}(\pi_{s}\xi)\right|^{p}d\mu,$ (11)
-we see that $m(|f^{\prime}(\pi_{s}\xi|^{p})=0$ if and only if $i(|f^{\prime}(\pi_{s}\xi)|)^{p}=0$ for $\mu$-a.e.
-$s\in S^{w}$.
-Lastly, we observe that $i(|f^{\prime}(\pi_{s}\xi)|)^{p}=0$ for $\mu$-a.e.
-$s\in S^{w}$ if and only if $i(|f^{\prime}(\pi_{s}\xi)|)^{2}=0$ for $\mu$-a.e.
-$s\in S^{w}$.
-∎
-Our next result can be seen as a generalization of [21, Theorem 16.34] (see also [31] and [22, Theorem 2.5]), and also relates the almost weakly stable vectors to the vectors from $\mathcal{H}_{w}$ in the compact-weak mixing decomposition.
+Let $S^{w}$ denote the weakly almost periodic compactification of $S$, and we let $i,M,$ and $\mu$ be as in Section 2.1. Let $f^{\prime}\in E^{\prime}$ be arbitrary. Since
+$m\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)=M\left(i\left(\left|f^{\prime}(\pi_{s}\xi)\right|^{p}\right)\right)=\int_{K(S^{w})}\left|f^{\prime}(\pi_{s}\xi)\right|^{p}d\mu,$ (10)
+we see that $m(|f^{\prime}(\pi_{s}\xi|^{p})=0$ if and only if $f^{\prime}(\pi_{s}\xi)=0$ for $\mu$-a.e.
+$s\in K(S^{w})$, and in fact for every $s\in K(S^{w})$ since this is a continuous function.
+Since $f^{\prime}$ is arbitrary, this implies $\pi_{s}\xi=0$ for $s\in K(S^{w})$.
+It is easy to see that $\pi$ extends to a representation of $S^{w}$ with the image in the weak closure of $\pi(S)$; by definition given in (1), we have $\xi\in E_{\text{aws}}$. ∎
+Our next result can be seen as a generalization of [29], [20, Theorem 16.34], and [21, Theorem 2.5].
+Furthermore, it relates the almost weakly stable vectors to the vectors from $\mathcal{H}_{w}$ in the compact-weak mixing decomposition.
 Theorem 4.2.
@@
   4. (iv)
-    $d_{*}-\displaystyle\lim_{g}\pi_{g}\xi=0\text{ weakly}$.
+    $d_{*}-\displaystyle\lim_{s}\pi_{s}\xi=0\text{ weakly}$.
 Moreover, if $S$ is locally compact, $\lambda$ is a Radon measure, and $(S,\lambda)$ possesses a left-Følner net $\mathcal{F}=(F_{i})_{i\in I}$, then (i)-(vii) are equivalent.
@@
   2. (vi)
-    $d_{\mathcal{F}}-\displaystyle\lim_{g}\pi_{g}\xi=0\text{ weakly}$.
+    $d_{\mathcal{F}}-\displaystyle\lim_{s}\pi_{s}\xi=0\text{ weakly}$.
   3. (vii)
-    $\underline{d}-\displaystyle\lim_{g}\pi_{g}\xi=0\text{ weakly}$.
+    $\underline{d}-\displaystyle\lim_{s}\pi_{s}\xi=0\text{ weakly}$.
 If we further assume $\mathcal{F}=(F_{n})_{n=1}^{\infty}$ is a Følner sequence, then (i)-(viii) are equivalent.
@@
     For some/each $p>0$, $\lim_{n\rightarrow\infty}\sup_{\|x^{\prime}\|\leq 1}\frac{1}{\lambda(F_{n})}\int_{F_{n}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}d\lambda(s)=0$.
-If we assume even further that $E^{\prime}$ is separable and $\lim_{n\rightarrow\infty}\lambda(F_{n})=\infty$, then (i)-(ix) are equivalent.
+If we assume even further that $E^{\prime}$ is (norm) separable and $\lim_{n\rightarrow\infty}\lambda(F_{n})=\infty$, then (i)-(ix) are equivalent.
   1. (ix)
-    There exists a set $A\subseteq S$ with $d_{\mathcal{F}}(A)=1$ such that $\lim_{\underset{s\in A}{s\to\infty}}T_{s}\xi=0$ weakly.
+    There exists a set $A\subseteq S$ with $d_{\mathcal{F}}(A)=1$ such that $\lim_{\underset{s\in A}{s\to\infty}}\pi_{s}\xi=0$ weakly.
 Proof.
 As discussed in Section 2.1, $m$ restricts to a bi-invariant mean on $W(S)$. We now use Lemma 4.1 to see that for some/each $p>0$ we have
-$E_{\text{aws}}$ $=\left\{\xi\in E\ |\ m\left(s\mapsto\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}\right)=0\ \forall\ x^{\prime}\in E^{\prime}\right\},$
+$E_{\text{aws}}$ $=\left\{\xi\in E\ :\ m\left(s\mapsto\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}\right)=0\ \forall\ x^{\prime}\in E^{\prime}\right\},$
 hence (i) and (ii) are equivalent.
@@
 $0=m\left(s\mapsto|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}\right)\geq\epsilon m\left(\mathbbm{1}_{U^{c}}(\pi_{s}\xi)\right)\text{, hence }m(\mathbbm{1}_{U}(\pi_{s}\xi))=1-m\left(\mathbbm{1}_{U^{c}}(\pi_{s}\xi)\right)=1.$ (11)
-To see that (iii) implies (ii), we observe that
+To see that (iii) implies (ii), we observe that for every $x^{\prime}\in E^{\prime}$ and $\epsilon>0$
 $m\left(s\mapsto|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}\right)\leq\epsilon m\left(\mathbbm{1}_{U}(\pi_{s}\xi)\right)+m\left(\mathbbm{1}_{U^{c}}(\pi_{s}\xi)\right)=\epsilon.$ (12)
-It is clear that (iv) implies (iii).
-To see that (iii) implies (iv), we recall that (iii) is equivalent to (i), and (i) is independent of the choice of $m$.
-In particular, if (iii) holds for a particular left-invariant mean, then (iii) holds for all left-invariant means, which yields (iv).
+It is clear that (iv) implies (iii). To see that (iii) implies (iv), it suffices to recall that $d_{*}\leq m$ for any left-invariant mean.
 Now let us assume that $(S,\lambda)$ admits a left-Følner net $\mathcal{F}=(F_{i})_{i\in I}$.
-We observe that $g(s):=|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}$ is in $W(S)$.
-For each $\in\mathcal{M}(\mathcal{F}), mm$ restricts to the unique bi-invariant mean on $W(s)$, so the value of $m(g)$ is independent of the choice of $m$, hence
-$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}d\lambda(s)=m(g)\text{ for all }m\in\mathcal{M}(\mathcal{F}).$ (14)
+We observe that $\phi(s):=|\langle\pi_{s}\xi,x^{\prime}\rangle|^{p}$ is in $W(S)$.
+For each $\in\mathcal{M}(\mathcal{F}), mm$ restricts to the unique bi-invariant mean on $W(S)$, so the value of $m(\phi)$ is independent of the choice of $m$, hence
+$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|^{p}d\lambda(s)=m(\phi)\text{ for all }m\in\mathcal{M}(\mathcal{F}).$ (13)
 In particular, the limit in the left hand side of (13) exists.
-Recalling that if (ii) holds for a particular left-invariant mean then it holds for all left-invariant means, we see that (ii) is equivalent to (v).
+We recall that (ii) is equivalent to (i), and (i) is independent of the choice of $m$.
+In particular, if (ii) holds for a particular left-invariant mean, then (ii) holds for all left-invariant means, which shows that (ii) is equivalent to (v).
 Similarly, we see that (iii) is equivalent to (vi).
-It is clear that (vii) implies (vi).
-To see that (vi) implies (vii), it suffices to observe that (vi) is equivalent to (i), so if (vi) holds for a particular left-Følner sequence $\mathcal{F}$, then (vi) holds for every left-Følner sequence $\mathcal{F}$.
+It is clear that (vii) implies (vi). To see that (vi) implies (vii), first note that, as above, the limit
+$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left\langle\pi_{s}\xi,x^{\prime}\right\rangle d\lambda(s)=d_{\mathcal{F}}-\displaystyle\lim_{s}\left\langle\pi_{s}\xi,x^{\prime}\right\rangle$
+exists for any $\xi\in E$ and $x^{\prime}\in E^{\prime}$, so that $\underline{d}(\{s\in S:\pi_{s}\in U\})=1$ implies $d_{\mathcal{F}}(\{s\in S:\pi_{s}\in U\})=1$ for every weak neighbourhood of zero $U$ in $E$, and for any left-Følner net $\mathcal{F}$.
 Now let us assume that $\mathcal{F}=(F_{n})_{n=1}^{\infty}$ is a Følner sequence.
 It is now clear that (viii) implies (v), so we will proceed to show that (v) implies (viii).
-Since $(\pi_{s})_{s\in S}$ is uniformly bounded, we suppose without loss of generality that $S$ is a monoid and that $(\pi_{s})_{s\in S}$ consists of contractions.
+Since $(\pi_{s})_{s\in S}$ is uniformly bounded, we suppose without loss of generality (as mentioned in Section 2.4) that $S$ is a monoid and that $(\pi_{s})_{s\in S}$ consists of contractions.
 Then the adjoint action $(\pi_{s}^{*})_{s\in S}$ leaves the ($\sigma^{*}$-compact) dual unit ball $D^{\prime}\subseteq E^{\prime}$ invariant.
@@
 $\lim_{n\rightarrow\infty}\frac{1}{\lambda(F_{n})}\int_{F_{n}}T_{s}f_{\xi}(x^{\prime})d\lambda(s)=\lim_{n\rightarrow\infty}\frac{1}{\lambda(F_{n})}\int_{F_{n}}f_{\xi}(\pi_{s}^{*}x^{\prime})d\lambda(s)=0$ (14)
-holds pointwise. Then, by the dominated convergence theorem the convergence of 15 holds weakly, i.e.,
+holds pointwise. Then, by the dominated convergence theorem the convergence of 14 holds weakly, i.e.,
 $\lim_{n\rightarrow\infty}\frac{1}{\lambda(F_{n})}\int_{F_{n}}\langle T_{s}f_{\xi},\mu\rangle d\lambda(s)=0$ (15)
 holds for any $\mu\in\left(C(D^{\prime})\right)^{\prime}$.
-By Eberlein’s generalized mean ergodic theorem, weak convergence implies the convergence of 16 in norm (see, e.g., Chapter 2, Theorem 1.5 of [34]).
+By Eberlein’s generalized mean ergodic theorem, weak convergence implies the convergence of 15 in norm (see, e.g., Chapter 2, Theorem 1.5 of [32]).
 Hence
@@
 Since $d_{\mathcal{F}}(C)=0$ for every compact set $C$, we see that (ix) implies (vi), so it remains to show that (vi) implies (ix).
-Since $E^{\prime}$ is separable, it is first countable, so let $\{U_{n}\}_{n=1}^{\infty}$ be a neighborhood basis for $0\in E^{\prime}$.
-For each $n\in\mathbb{N}$, let $A_{n}=\{s\in S|\pi_{s}\xi\in U_{n}\}$ and let $m_{n}$ be such that $\lambda(A_{n}\cap F_{m})>\frac{n-1}{n}\lambda(F_{m})$ for all $m\geq m_{n}$.
-Assuming without loss of generality that $m_{n+1}>m_{n}$, we see that it suffices to take $A=\bigcup_{n=1}^{\infty}\bigcup_{m=m_{n}+1}^{m_{n+1}}(A_{n}\cap F_{m})$.
-∎
+As in the previous case, we can suppose that every $\pi_{s}$ is a contraction.
+Since $E^{\prime}$ is separable, the weak topology on $E$ is first countable when restricted to the unit ball $E_{1}$, so let $\{U_{n}\}_{n=1}^{\infty}$ be a weak neighborhood basis for 0 in $E_{1}$.
+We can suppose it is decreasing, and to fix notations, set $U_{n}=V_{n}\cap E_{1}$ with $V_{n}$ weakly open in $E$.
+For each $n\in\mathbb{N}$, set $A_{n}=\{s\in S\ :\ \pi_{s}\xi\in U_{n}\}$; since every $\pi_{s}$ is contractive, we have also $A_{n}=\{s\in S\ :\ \pi_{s}\xi\in V_{n}\}$, so by (vi), $d_{\mathcal{F}}(A_{n})=1$.
+Let $m_{n}$ be such that $\lambda(A_{n}\cap F_{m})>\frac{n-1}{n}\lambda(F_{m})$ for all $m\geq m_{n}$.
+Assuming without loss of generality that $m_{n+1}>m_{n}$, we see that for $A:=\bigcup_{n=1}^{\infty}\bigcup_{m=m_{n}+1}^{m_{n+1}}(A_{n}\cap F_{m})$ we have $d_{\mathcal{F}}(A)=1$ and
+$\{s\in A:\pi_{s}\xi\notin U_{n}\}=A\setminus A_{n}\subset\bigcup_{k<n}\;\bigcup_{m=m_{k}+1}^{m_{k+1}}(A_{k}\cap F_{m})\subseteq\bigcup_{m=1}^{m_{n}}F_{m}.$
 Example 4.3.
-We now recall [32, Example 4.3] to give an example showing that Lemma 4.1 does not necessarily hold if $W(S)$ only admits a left-invariant mean and not a bi-invariant mean.
+We now recall [30, Example 4.3] to show that Lemma 4.1 does not necessarily hold if $W(S)$ only admits a left-invariant mean and not a bi-invariant mean.
 This example will also show that Theorem 4.2 does not necessarily hold if $S$ is a left-amenable subsemigroup of an amenable group that is not right amenable.
-We consider the semigroup $(S,\cdot)$ given by $S=\{(m,n)\in\mathbb{Z}^{2}\ |\ m\geq 0,n\geq 1\}$ and $(m_{1},n_{1})\cdot(m_{2},n_{2})=(2^{n_{2}}m_{1}+m_{2},n_{1}+n_{2})$.
-$S$ is cancellative and left-amenable, hence it satisfies the strong Følner condition and possesses a left-Følner sequence $\mathcal{F}=(F_{i})_{i=1}^{\infty}$.
+We consider the semigroup $(S,\cdot)$ given by $S=\{(m,n)\in\mathbb{Z}^{2}\ :\ m\geq 0,n\geq 1\}$ with the multiplication
+$(m_{1},n_{1})\cdot(m_{2},n_{2})=(2^{n_{2}}m_{1}+m_{2},n_{1}+n_{2}).$
+As shown in [30], $S$ is cancellative and left-amenable, hence it satisfies the strong Følner condition, and since it is countable, it possesses a left-Følner sequence $\mathcal{F}=(F_{i})_{i=1}^{\infty}$.
+Furthermore, by replacing $F_{n}$ with $F_{n}\cdot(0,1)$ which form still a left Følner sequence, we may assume without loss of generality that $F_{n}\subseteq S\cdot(0,1)$ for all $n$.
+We observe that
+$S\cdot(0,1)=\{(m,n)\in S\ :\ m\in 2\mathbb{Z},\;n\geq 2\},\quad S\cdot(1,1)=\{(m,n)\in S\ :\ m\notin 2\mathbb{Z},\;n\geq 2\}.$
 Since $S\cdot(0,1)\cap S\cdot(1,1)=\emptyset$, we see that $S$ is not right amenable.
-Furthermore, by replacing $F_{n}$ with $F_{n}\cdot(0,1)$, we may assume without loss of generality that $F_{n}\subseteq S\cdot(0,1)$ for all $n$.
-We observe that $S=S\cdot(0,1)\cup S\cdot(1,1)\cup\mathbb{Z}_{\geq 0}\times\{0\}$.
-Now consider the function $f:S\rightarrow\{0,1\}$ given by $f(s)=1$ if and only if $s\in S\cdot(0,1)$.
-We see that $\{R_{s}f\}_{s\in S}=\{f,0,1,\mathbbm{1}_{A}-f\}$, where $A=S\cdot(0,1)\cup S\cdot(1,1)$, hence it is relatively weakly compact and we see that $f\in W(S)$.
-Now let $E=W(S)$, $\xi=f$, $\pi=R$, and $x^{\prime}$ denote evaluation at $s_{0}$ for some $s_{0}\in S$.
-We see that $f\in E_{\text{aws}}$ because $R_{(1,1)}f=0$.
-However, if $m\in\mathcal{M}(\mathcal{F})$, then
-$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|d\lambda(s)=m(s\mapsto\left(R_{s}f\right)(s_{0}))=m(1)=1.$ (17)
+Now let $\xi:S\rightarrow\{0,1\}$ be the characteristic function of $S\cdot(0,1)$.
+We note that for $t\in S$ and $s=(m,n)\in S$, we have $ts\in S\cdot(0,1)$ if and only if $m\in 2\mathbb{Z}$.
+We see that $\{R_{s}\xi\}_{s\in S}=\{0,1\}$, hence it is relatively weakly compact and $\xi\in W(S)$.
+Now set $E=W(S)$, $\pi=R$, and let $x^{\prime}$ denote the evaluation at $s_{0}$ for some $s_{0}\in S$.
+We see that $\xi\in E_{\text{aws}}$ because $R_{(1,1)}\xi=0$.
+However, for every $s\in F_{i}\subset S\cdot(0,1)$ we have $R_{s}\xi=1$, thus if $m\in\mathcal{M}(\mathcal{F})$, then
+$\lim_{i}\frac{1}{\lambda(F_{i})}\int_{F_{i}}\left|\left\langle\pi_{s}\xi,x^{\prime}\right\rangle\right|d\lambda(s)=m(s\mapsto\left(R_{s}\xi\right)(s_{0}))=m(1)=1.$ (16)
 5 The JdLG decomposition via ultrafilters
 Let $E=\mathcal{H}$ be a Hilbert space, $U$ a unitary operator on $\mathcal{H}$, and $\mathcal{H}=\mathcal{H}_{c}\oplus\mathcal{H}_{w}$ the compact-weak mixing decomposition.
-There are various results in the literature, such as [8, Theorem 2.2], [10, Theorem 2.25], [35, Theorem 4], and [12, Section 6], that (allow one to) construct the orthogonal projection $P:\mathcal{H}\rightarrow\mathcal{H}_{c}$ as an ultrafilter limit of powers of $U$.
+There are various results in the literature, such as [8, Theorem 2.2], [9, Theorem 2.25], [34, Theorem 4], and [11, Section 6], that (allow one to) construct the orthogonal projection $P:\mathcal{H}\rightarrow\mathcal{H}_{c}$ as an ultrafilter limit of powers of $U$.
 Theorem 5.2 is motivated by these previously mentioned results and attempts to generalize them to JdLG-admissible representations on a Banach space $E$.
@@
 In particular, $h$ sends idempotents to idempotents.
-Since $\mathscr{S}_{r}=K(\mathscr{S})$ is a compact topological group, so it contains a unique idempotent $Q$.
+Since $\mathscr{S}_{r}=K(\mathscr{S})$ is a compact topological group, it contains a unique idempotent $Q$.
 Furthermore, we have already seen that $Q$ acts as the identity on $E_{r}$, so the desired result follows.
-We now proceed to prove (ii). Theorem 4.2 tells us that for any weakly open neighborhood $U$ of 0, we have
-$d^{*}\left(\left\{s\in S\ |\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\right)=0\text{, hence }\left\{s\in S\ |\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\notin q.$ (18)
-Since $q$ is an ultrafilter, we see that $\left\{s\in S\ |\ \pi_{s}\xi_{\text{aws}}\in U\right\}\in q$. Since $U$ was arbitrary, we see that $q-\lim_{s}\pi_{s}\xi_{\text{aws}}=0$. ∎
+We now proceed to prove (ii). Theorem 4.2 tells us that for any weakly open neighborhood $U$ of 0 and every $\xi_{\text{aws}}\in E_{\text{aws}}$, we have
+$d^{*}\left(\left\{s\in S\ :\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\right)=0\text{, hence }\left\{s\in S\ :\ \pi_{s}\xi_{\text{aws}}\notin U\right\}\notin q.$ (17)
+Since $q$ is an ultrafilter, we see that $\left\{s\in S\ :\ \pi_{s}\xi_{\text{aws}}\in U\right\}\in q$. Since $U$ was arbitrary, we see that $q-\lim_{s}\pi_{s}\xi_{\text{aws}}=0$. ∎
 Theorem 5.2.
@@
 It follows that $K(\mathscr{S})=\phi(K(\beta S))$.
-Since $\pi$ is a JdLG-admissible representation, we see that $K(\mathscr{S})$ is a compact topological group with a unique idempotent $Q$, which is the projection onto $E_{r}$.
+Since $\pi$ is a JdLG-admissible representation, we see that $K(\mathscr{S})$ is a compact topological group with a unique idempotent $Q$, which is the projection onto $E_{r}$ along $E_{\text{aws}}$.
 Since $K(\mathscr{S})$ is closed, we see that $\phi\left(\overline{K(\beta S)}\right)=K(\mathscr{S})$.